A Complete Bibliography of Publications in the International Journal of Statistics and

Probability

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

02 June 2023 Version 1.12

Title word cross-reference

1 [311]. 2^{m} [222]. 2 × 2 [272]. 9 [546]. A [526]. D [474, 119, 476, 577]. G [455, 488, 307, 343, 406, 302]. Γ [6]. k [299]. LR [20]. LXB = 0 [129]. m [109]. L_{2} [166]. R^{∗}({1}|Ω_{ℓ}) [222]. N < ν_{ℓ}(m) [222]. n < v [157]. P [589]. P (Y < X) [142]. P = 0 [372]. P = 1 [372]. Q [531]. R [539]. RS [48]. T [605]. T − X [65]. T^{2} [317, 465].

-Component [546]. -distribution [605]. -learning [531]. -Memory [109]. -Optimal [119, 476, 526, 577, 474]. -squared [539]. -structure [166]. -value [589].

1 [340, 413, 470, 341]. 19 [499, 510, 496, 589, 516].

2 [342, 200, 610].

3 [211].

4 [301].

5 [314].

6 [331].

8 [556].

9-Component [547].

Aalen [277]. Ability [515]. Absenteeism [556]. Absolute [22, 192]. Absorbing [225]. Academic [122]. Accelerated [99, 315, 334]. Acceleration [338]. Accounting [140]. Accrual [17]. Accuracy [468, 309, 173]. Acknowledgements [100, 112, 120, 138, 152, 170, 180, 191, 200, 211, 228, 235, 247, 262, 275, 293, 301, 314, 331, 341, 351, 361, 373, 384, 399, 412, 427, 436, 445, 454, 464, 472, 478, 486, 493, 503, 514, 522, 534, 545, 560, 569, 576, 582, 588, 594, 600, 606, 613, 619]. Across [566, 246, 443]. Actually [581]. Acute [5]. Adaboost [375]. Adapted [300]. Adaptive [352, 391, 31, 187, 233, 310, 172]. Additive [307, 198, 184]. Adjusted [294, 121]. Adjusting [81]. Adolescents [297]. Adults [115]. Advanced [443]. Advantage [161]. Affine [225]. after [233]. Age [224, 581, 463]. Aged [595]. Aggregate [55]. Agreement [441]. Agricultural [557]. AIDS [489]. Air [587]. Akuapem [379]. Alabama [378]. Algebras [461]. Algorithm [485, 118, 148, 189, 517, 71]. Algorithmic [475]. Algorithms [559, 249, 448, 363]. Allocating [31]. Allocation [487, 518, 367]. Allocations [98]. Allowing [350]. Alpha [578, 584, 511]. Alternative [214, 329]. Alternatives [347]. Altman [479]. Among [533, 94, 595, 70, 463, 297]. Analogy [431]. Analysed [167]. Analyses [608]. Analysis [245, 52, 595, 500, 367, 377, 209, 139, 437, 36, 103, 441, 513, 602, 415, 490, 151, 502, 481, 479, 386, 43, 320, 438, 537, 452, 116, 426, 92, 466, 387, 618, 280, 26, 286, 587, 593]. Analytical [543]. Analyzing [323, 556, 496]. ANN [599]. Annual [480]. Annuities [447]. Answers [309]. Antenatal [354]. Anthropometric [92]. Anti [292, 363]. Anti-Epileptic [363]. Anti-infective [292]. API [88]. Applicable [306]. Application [94, 573, 253, 404, 102, 147, 328, 455, 330, 35, 300, 584, 601, 322, 342, 513, 415, 495, 556, 453, 526, 389, 343, 6, 589, 294, 444, 466, 554, 234, 363, 525, 305, 199, 354, 14, 178]. Applications [383, 179, 585, 368, 154, 65, 459, 538, 260, 407, 488, 181, 370, 578, 337, 45, 311, 385, 532, 550, 406, 598, 511, 553, 536, 145, 357, 132, 2, 286]. Applied [273, 603, 428]. Applying [310]. Approach [267, 487, 518, 102, 238, 11, 59, 513, 559, 497, 44, 479, 531, 240, 475, 548, 20, 591, 57, 136, 291, 168, 155]. Approaches [77, 424, 390]. Approximate [269, 104]. approximating [471]. Approximation [592, 98, 428, 366]. Approximations [18, 145, 121]. Archimedean [542, 562, 471]. Area [102, 248, 386]. Areas [35, 378]. ARFIMA [44]. ARHMM [448]. ARIMA [456, 604]. Arising [457]. Arithmetic [4]. arm [402]. Arms [81]. ART [305]. Artificial [599, 88, 196]. Aspect [219]. Aspects [423, 203]. Assess [557, 67]. Assessing [350, 566, 441, 246, 447, 150, 78, 115]. Assessment [236, 232, 326]. Assessments [288]. Asset [487, 518]. Associated [84, 559, 297]. Association [139, 5, 593]. Associations [269]. Asthma [587]. Asymmetric [325, 10, 316]. Asymmetrical [521]. Asymptotic [316, 188, 393, 101, 155]. Asymptotically [206, 47]. Atmospheric [19]. Attachment [97, 128]. Attitude [544]. Attribute [441]. Attrition [323]. Auto [480, 591, 448]. Auto-Regressive [480, 448]. Autocorrelated [554, 196]. Autocorrelation [20]. Automaton [283]. Autoregressive [599, 263, 456, 296]. Auxiliary [265, 206]. Average [77, 411, 599, 61, 480, 456]. Average-ARIMA [456]. Averages [393].

B [410]. B-spline [410]. Background [69]. Backward [448]. Bad [614]. Balanced [360, 222]. Bank [379, 204]. Banks [204]. Bar [467]. Bar-Lev [467]. Based [15, 21, 22, 106, 572, 468, 494, 252, 565, 391, 238, 469, 122, 347, 289, 428, 372, 564, 345, 496, 295, 53, 131, 546, 547, 116, 92, 364, 176, 48, 283, 291, 155, 410, 375, 611, 266, 313, 20]. Basic [97]. Basin [602]. Basis [607, 528, 264]. Basketball [131, 161]. Bathtub [538, 142, 176]. Baye [94]. Bayes [468, 491, 55, 371, 537, 288, 315]. Bayes-Estimation [491]. Bayesian [15, 487, 518, 102, 607, 258, 333, 174, 411, 35, 273, 59, 513, 590, 477, 527, 567, 442, 416, 452, 232, 85, 75, 205, 140, 617, 611]. Bayesian-type [174]. be [318]. Behaved [431]. Behavior [114]. Behaviour [119]. Benefits [447]. Bent [19]. Bent-cable [19]. Bernoulli [110, 479]. Best [151]. Beta [95, 574, 389, 343, 53, 302, 38, 14]. Beta- [302]. Beta-Binomial [53]. Better [208]. Between [287, 395, 268, 73, 92, 431, 591, 219, 5, 50]. Bias [566, 38]. Biased [374]. Biases [579]. Bilateral [348]. Bilinear [166]. Binary [162, 8, 274, 66, 125, 57, 390, 135]. Binomial [149, 592, 353, 366, 53, 132, 612]. Bioequivalence [402]. Biomarkers [240]. Birth [548, 581]. Birth-and-Death [548]. Bitcoin [502]. Bivariate [451, 336, 504, 609, 590, 605, 60, 125, 471, 116, 63, 348, 612, 306]. Black [523]. Black-Litterman [523]. Bladder [14]. Bland [479]. Blending [526]. BLUEs [512]. BMI [297]. Boltzmann [543, 549]. Bolus [105]. Boosted [414]. Bootstrap [267, 565]. Both [231]. Bound [429]. Boundaries [253, 237]. Boundary [434, 225, 376]. Bounds [563, 366]. Bradley [7]. Brain [320]. Breaks [76]. Breast [94, 136]. Bridge [73, 41]. Bridges [261]. Brief [284]. Brownian [225, 32, 261]. BSDEs [540]. Buffered [491]. Burkina [510]. Burr [573, 380, 149, 89, 407, 584, 550, 132, 178]. Business cable [19]. Calculation [199, 402]. Calibration [48]. Can [208, 318]. Canada [496]. Cancer [94, 273, 69, 495, 240, 136, 193, 14]. Canonical [116, 26]. Capita [551]. Carbon [392]. Care [354]. Carlo [432, 67]. Case [379, 204, 528, 365, 568, 104]. Case-Crossover [568]. Cases [290, 597, 516]. Categorical [371, 329]. Categories [216, 28]. Cauchy [160]. Causal [517]. Cavitation [425]. CD4 [305]. Cellular [365, 283]. Censored [164, 146, 106, 11, 469, 202, 490, 176, 611]. Censoring [494, 421, 72, 561, 334, 108, 104]. Center [171]. centered [171]. Central [72, 171, 382, 144, 567]. Certain [335, 355, 405]. CFA [491]. CFC [19]. Chain [245, 579, 430, 548, 480, 80]. Chains [442]. Change [83, 103, 505, 537, 444, 193, 233]. Change-point [233]. Changed [581]. Changepoint [153]. Changepoints [327]. Changes [23]. Changing [119]. Chaos [236]. Characteristic [524]. Characteristics [404]. Characterization [374, 333, 126, 337, 254]. Characterizations [163, 149, 312, 156, 260, 407, 110, 160, 335, 355, 405, 439, 482, 289, 256, 418, 570, 284, 512]. Chart [22]. Chi [82]. Chi-Square [82]. child [358]. Childbearing [463]. China [358]. Cholera [430]. Cholesky [74]. Circularly [66]. Circulating [136]. CKLS [127]. Claim [199]. Claims [437, 618]. Class [572, 467, 377, 360, 69, 385, 406, 38, 615]. Classes [54]. Classification [468, 515]. Classifier [408]. Classifiers [468]. Classroom [267]. Climate [444]. Clinical [17, 218, 402, 261, 231]. Closing [480]. Closure [54]. Cluster [217]. Clustered [381, 271]. Clustered-Task [271]. Clustering [109]. Clusters [491]. Co [475, 443, 39, 387]. Co-Evolution [475]. Co-movements [443]. Co-ordination [39]. Co/Multimorbidity [387]. Coach [131]. Coating [525]. Coefficient [460, 63, 280]. Coefficients [166, 347, 540]. Coherent [165, 431]. Cohort [587]. Collapsed [215]. Collection [485]. Collective [458]. College [131, 161]. Collinearity [51]. Collisions [41]. Colorectal [69]. Colored [515]. Com [385]. Com-Poisson [385]. Combined [485, 305]. Combining [608]. Comments [319]. Commodity [328, 295]. Common [299]. Companies [379]. Comparative [533, 212, 396, 148]. Compare [376]. Comparing [111, 107, 220, 480]. Comparison [164, 146, 183, 209, 338, 378, 372, 53, 484, 408, 471, 596]. Comparisons [7]. Compartment [105]. Competing [277]. Competitor [40]. Complementary [273, 364]. Complete [469, 93]. Completely [156]. Component [367, 546, 547, 426, 466]. Components [184]. Composite [415, 171, 382, 144]. Compound [435, 542, 562]. Comprehensive [245]. Computation [432, 225]. Computational [423]. Concept [583]. Concomitants [63]. Condition [367]. Conditional [163, 36, 337, 72, 216, 266, 433, 187, 113]. Conditions [59, 222, 446]. Confidence [267, 344, 565, 521, 579, 465, 505]. Configural [234]. Confirmatory [390]. Congeneric [229]. Connections [365, 73]. Considering [231]. Consistency [519, 192]. Constant [106, 17]. Constraint [173]. Constraints [177]. Construct [234]. Construct-focused [234]. Construction [382, 39]. Consumption [531, 115]. Contagion [499].

Contaminated [84]. Contamination [188]. Context [607]. Contextual [446]. Contingency [216, 359, 605, 422, 4, 49, 215, 272]. Continuous [77, 166, 439, 430]. Continuous-time [166]. Contoured [24]. Contraception [492]. Contradictions [221]. Contrasts [291]. Control [22, 568]. Controlled [378]. Convergence [194, 277, 16, 93, 185]. Convergent [118, 148, 189]. Convex [227, 276, 498, 519, 564, 192]. Convexity [303]. Convexity/Monotonicity/Positivity [303]. Convolution [238, 461]. Convolutional [414]. Conway [504]. Copula [609, 590, 43, 266, 220, 471, 546, 547]. Copula-based [266]. Copulas [336, 398, 36, 542, 562, 279, 408, 596]. Corona [510]. Corona-Virus [510]. Corrected [434]. Correcting [311]. Correction [38, 223]. Correlated [608, 223]. Correlation [523, 607, 271, 280, 26]. Correlation-Preserving [607]. Correlations [359, 305]. Correspond [318]. Cotton [386]. Count [609, 311, 558, 305, 354]. Counting [397, 66]. Countries [557]. Counts [452]. Courses [362]. Court [161]. CoV [610]. Covariance [25, 229, 75, 129, 512, 26]. Covariate [586]. Covariates [282]. COVID [499, 510, 496, 589, 516, 618]. COVID-19 [499, 510, 496, 589, 516]. Cox [109, 616, 145]. Cox-Type [109]. Coxian [46]. Crack [324]. Cramér [91, 188]. Crash [369, 243, 378]. Crashes [204]. Credit [3]. Criteria [520]. Criterion [186]. Cross [528]. Cross-Sectional [528]. Crossing [504]. Crossover [568]. Crude [591]. Cube [171]. Cubic [407]. Cultivated [386]. Cumulative [289, 567]. Cure [213, 590, 101]. Current [565]. Curve [248]. Customer [323]. CUSUM [537, 159, 233]. Cutoff [589]. Cycles [443].

Dagum [270]. Daily [411]. Data [164, 146, 380, 183, 106, 538, 575, 27, 565, 608, 162, 181, 381, 330, 230, 528, 609, 584, 601, 273, 342, 11, 45, 202, 415, 490, 175, 151, 497, 60, 453, 562, 153, 481, 311, 274, 371, 558, 6, 388, 111, 329, 416, 589, 294, 47, 509, 92, 554, 169, 132, 176, 57, 234, 390, 85, 393, 473, 354, 306, 611, 178, 578]. Database [27]. Datas [598]. Death [548]. Deaths [516]. Decision [520, 111, 414, 168, 137]. Decisions [208]. Decline [19]. Decomposability [34]. Decomposition [74, 359, 527, 426, 470, 215]. Deep [616]. Definite [512]. Degradability [603]. Degree [128, 476, 526]. Demand [562]. Demographic [595]. Demystifying [583]. Density [84, 207, 195, 428, 349, 121, 306]. Departure [422, 7]. Dependence [243, 279, 90, 266]. Dependency [398]. Dependent [42, 300, 236, 210, 251, 68, 135, 393, 108]. Derivation [420, 426]. Derivative [328, 83, 103]. Design [474, 214, 189, 476, 526, 452, 144, 568]. Designs [391, 360, 222, 119, 171, 31, 577, 382]. Detect [401]. Detectability [489]. Detecting [268, 71]. Detection [325, 23, 83, 236, 490, 159]. Determinants [245, 338]. Determination [237]. Determining [297]. Deterministic [239]. Developing [483]. Development [407, 12, 167]. Developments [197]. Deviance [186]. Deviation [22, 128, 130]. Deviations [97, 192]. DHGF [131]. Diabetes [483, 290, 342]. Diabetic [11, 59]. Diagnostic [546, 547]. Diagnostics [317]. Diagonal [216]. Dichotomy [219]. Dickey [376]. Diet [593]. Diet-Disease [593]. Difference [268, 85, 612]. Different [10, 474, 270, 86, 321]. Differential [308, 136]. Difficulty [446, 264]. Diffusion [253, 300]. Dimensional [60, 68, 234]. Direction [17]. Directional [258]. Disaggregation [183]. Disclosure [429]. Discovery [83]. Discrete [325, 592, 54, 352, 77, 405, 439, 556, 311, 564, 570, 561]. Discrete-Time [325]. Discriminant [408]. Discriminants [596]. Discriminate [287]. Discriminating [70]. Discrimination [515]. Disease [290, 601, 69, 495, 510, 442, 593]. Dispatch [257]. dispersed [274]. Dispersion [203]. Distance [347, 408, 68, 20]. Distance-Based [347]. Distinguish [597]. Distribution [380, 383, 15, 179, 312, 573, 106, 89, 592, 212, 423, 585, 417, 494, 154, 396, 226, 267, 102, 260, 147, 324, 374, 407, 504, 298, 344, 190, 281, 278, 333, 270, 435, 330, 370, 230, 578, 584, 110, 126, 322, 259, 45, 160, 469, 495, 127, 453, 449, 477, 339, 400, 91, 254, 419, 543, 549, 558, 574, 389, 550, 53, 6, 570, 321, 561, 541, 302, 188, 159, 471, 198, 598, 511, 553, 536, 507, 294, 357, 552, 39, 24, 332, 132, 63, 142, 176, 615, 85, 48, 580, 269, 129, 250, 14, 611, 178]. Distribution [286, 605]. Distribution-preserving [129]. Distributionm [501]. Distributions [164, 146, 21, 163, 149, 54, 585, 156, 65, 459, 352, 398, 455, 488, 95, 98, 609, 337, 335, 355, 405, 439, 482, 462, 307, 428, 418, 385, 343, 532, 53, 406, 73, 145, 232, 284, 348, 612, 201, 46]. District [492]. Disturbances [196]. Divisibility [335, 355]. Does [173, 492, 221]. Domestic [544]. Dose [124]. dot [461]. Double [206, 215]. Drift [225]. Driven [32, 336]. Driving [446]. Drug [363, 525]. Due [490]. During [520, 305]. Dynamic [523, 239, 217, 131, 58]. Dynamics [425, 430].

E-Bayesian [611]. Each [199]. Eco [557]. Eco-efficiency [557]. Economic [183, 481, 438, 551]. Economies [443]. Edgeworth [555]. Education [544, 492]. Effect [544, 391, 224, 517, 586]. Effects [325, 595, 96, 232, 75, 242]. Efficacy [520, 125, 124]. Efficiencies [214]. Efficiency [161, 196, 101, 557]. Efficient [527, 29, 473]. EGTG [332]. Eigenvalues [318]. Electricity [562]. Elliptically [24]. Emerging [169]. Empirical [219, 42, 209, 98, 128, 130, 521, 481, 29, 535, 410]. Empty [264]. Empty-loaded [264]. Endpoint [402]. Endpoints [124]. Energy [279, 168]. Enhanced [288]. Enhancement [468, 375]. Enis [467]. Ensemble [468]. Entropy [237, 289]. Entry [505]. Environment [168]. Environmental [177]. Epidemic [430, 548]. Epidemics [489]. Epidemiological [209, 51]. Epileptic [363]. Epistatic [96]. Epistemological [39]. Equal [63]. Equality [258, 87, 512]. Equally [107]. Equation [533, 607, 308, 539, 242]. Equations [58]. Equivalence [117]. equivalent [28]. Era [96]. Error [477, 424, 408, 4, 356]. Errors [223]. Essential [431]. Estate [283]. Estimate [194, 84, 11, 113]. Estimated [311]. Estimates [244, 62]. Estimating [352, 265, 248, 411, 290, 299, 59, 18, 243, 498, 125, 96, 50, 135, 242]. Estimation [533, 21, 212, 157, 10, 253, 494, 252, 203, 491, 265, 565, 270, 421, 230, 309, 34, 513, 469, 559, 175, 207, 195, 394, 249, 477, 524, 400, 90, 197, 227, 276, 349, 321, 561, 517, 68, 507, 308, 552, 63, 187, 142, 173, 75, 580, 205, 201, 184, 473, 233, 104, 306, 26, 356, 2, 611]. Estimations [239, 336]. Estimator [555, 434, 72, 460, 484, 277, 134, 167, 193]. Estimators [533, 244, 206, 372, 192, 1]. Ethnicity [463]. Euclidean [408]. European [557]. Evaluating [115, 506, 264, 283]. Evaluation [177, 391, 521, 122, 131]. Event [239, 109, 135, 108]. Event-Dependent [108]. Events [346]. Evidence [443, 29]. Evolution [475]. Exact [123, 119, 402]. Examination [204]. Examining [463]. Example [12]. Examples [420]. Excel [448]. Excess [199, 354]. Exchange [234]. Exchangeability [346]. Exciting [109]. Exercise [115]. Exercising [447]. Existence [17, 222, 58]. Exists [188]. Expansion [555]. Expectancy [43]. Expectation [163, 433]. Expectations [337]. Expected [421, 462]. Experiment [476, 526, 515]. Explained [135]. Explaining [362]. Explicit [127, 512]. Exploration [544]. Exponent [236, 216]. Exponential [380, 312, 203, 467, 98, 230, 273, 225, 495, 457, 18, 256, 78, 321, 294, 101, 233, 332]. Exponential-Generalized [332]. Exponentiated [15, 494, 65, 278, 330, 370, 421, 482, 449, 570, 198, 282, 615, 14, 178]. Exponentiated-G [615]. Export [114]. Expressions [24]. Extended [89, 423, 417, 316, 298, 435, 578, 322, 256, 511, 553, 364, 7, 193, 49]. Extension [346, 281, 291]. Extra [117]. Extremal [36]. Extreme [147, 437, 256, 254, 444, 199].

Face [177, 171]. Face-centered [171]. Factor [244, 150, 371, 537, 577, 173, 390, 446, 375]. Factorial [222]. Factors [523, 483, 350, 378, 506, 297]. Failure [538, 338, 204, 99, 507, 142, 580, 108]. Falling [167]. Falls [167]. False [83]. Families [149, 154, 459, 467]. Family [585, 397, 65, 82, 328, 455, 488, 95, 307, 343, 532, 406, 233]. Faso [510]. Fatalities [246]. Fatality [224]. Fear [167]. Fertility [595, 463]. Field [540, 8]. Filtered [83, 103]. Finance [253]. Financial [379, 243, 2]. Finding [124]. Finite [434, 524, 8]. First [253, 581, 458]. First-passage [253]. Fisher [408, 296]. Fisheries [388]. Fistula [338]. Fit [78, 173, 40, 497, 450]. Fitness [97]. Fitting [562, 551]. Fixed [577]. Flaring [392]. Flexibility [168]. Flexible [383, 202, 425, 125, 302, 75]. Flood [575, 602]. Flow [183]. Flows [295]. focused [234]. Force [80]. Forecast [566, 510]. Forecasting [456, 604, 196, 2]. Forecasts [263]. Fork [123]. Fork-Join [123]. Form [229, 321]. Formulas [420]. Forward [448]. Four [439, 86, 526]. Fourier [241]. Fractional [222, 105, 540]. Frailty [252, 338]. Framework [33, 372, 456]. Frechet [536, 179, 585]. Free [57]. Frequency [602, 401]. Frequentist [208]. Frontier [557]. Fruits [526]. Fuller [376]. Function [592, 585, 397, 292, 194, 374, 225, 55, 477, 254, 408]. Functional [500, 60, 320, 37, 410]. Functions [156, 10, 467, 559, 590, 87, 227, 276, 303, 498, 564, 192, 420, 456]. Fundamental [93]. Future [17, 295]. Fuzzy [251, 168, 137].

G [532, 585, 584, 615]. Gains [431]. Game [442]. Games [8]. Gamma [82, 298, 202, 453, 348, 286]. Gap [34]. Gaps [91]. GARCH [325, 316, 266]. Gaussian [483, 546, 547, 68]. GDP [551]. Geeta [400]. General [572, 34, 484, 321, 456, 117, 129, 512]. Generalised [85]. Generalizations [541]. Generalized [15, 21, 163, 154, 396, 459, 147, 336, 298, 455, 278, 237, 482, 307, 151, 453, 37, 53, 406, 73, 294, 357, 80, 282, 540, 48, 155, 250, 286, 332]. Generate [150]. Generated [260, 202]. Generating [394]. Generation [604, 591]. Generator [418]. Genes [94]. Genetic [218, 345]. Genomic [96]. Geographical [246]. Geometric [346, 298, 130, 435, 110, 273, 482, 457, 506, 332, 612, 4, 280]. Geometric-Zero [482]. Geometrical [219]. Geostatistical [36]. Getting [264]. Ghana [114, 595, 379, 224, 604, 591]. Gibbs [185, 525]. Girsanov [127]. Given [246]. GLM [385]. Global [481, 551]. Good [614]. Goodness [40, 497, 78]. Goodness-of-fit [40, 497]. Google [88]. graded [527]. Gradient [213, 84]. Graph [78]. Graphs [98, 128, 130]. Greedy [485]. Group [575, 240]. Grouping [377]. Growth [287, 481]. Guaranteed [447]. Gumbel [573]. GWR [440].

Haghighi [449]. Half [396, 278, 160, 532]. Half-Logistic [396, 278]. Handling [575, 53, 424]. Hannan [30]. Harmonic [4]. Harris [598, 553]. Hatke [380]. Hazard [156, 72, 505, 176]. Health [595, 59, 618]. Healthcare [441]. Heart [601]. Heavy [324, 428]. Heavy-tailed [324, 428]. Hellinger [68]. Help [208]. Heterogeneity [338]. Heterogeneous [123]. Heteroscedasticity [474, 388, 424, 159]. Hidalgo [610]. Hidden [186, 448]. Hierarchical [596]. High [401, 151, 115]. High-frequency [401]. High-Intensity [115]. High-Speed [151]. Higher [223, 20, 492]. Hill [37]. HIV [305, 489]. HMM [295]. Hoeffding [563]. Home [161, 618]. Homogeneity [381, 364, 49]. Homogeneous [396, 80]. Hospital [320]. Hosts [475]. Hotelling [395, 317]. Household [5]. Housing [283]. Hybrid [239, 288]. Hydroelectric [591]. Hydropower [604]. Hypothesis [329, 61, 29, 440].

Identical [467]. Identifiability [296]. Identification [448, 240]. Identifying [189, 136, 305, 614]. Identity [429]. II [536, 585, 494, 421, 584, 561, 124, 334, 104, 611]. II-G [585]. III [407, 335, 339]. Illustrate [267]. Illustrative [420]. Imaginary [174]. Impact [369, 481, 67, 51]. Impacts [350, 492]. Impair [173]. Importance [12, 150]. Improved [429, 563, 1, 49]. Improvement [27]. Improving [441, 309]. Imputation [27]. Income [572, 447, 5]. Incomplete [572]. Incorporating [69, 617]. Increase [329]. Independence [359, 129, 272]. Independent [467, 491, 457, 366]. Index [64, 451, 77, 34, 224, 158]. Indicator [143, 135]. Indicators [304]. Indifference [221]. Individuals [492]. Induced [444]. Industry [245]. Inequalities [145, 79]. Inequality [572]. Infected [305]. Infection [610]. infective [292]. Inference [106, 174, 578, 44, 99, 442, 511, 553, 536, 313, 232, 315, 617, 410]. Inferences [370, 176, 586]. Inferential [62]. Inferring [94]. Infestation [567]. Infinite [335, 355]. Inflated [609, 38, 463, 201, 190, 175, 13]. Inflow [118, 148, 189]. Influence [554, 264, 358]. Informal [62, 313]. Information [97, 186, 296, 101, 333]. Ingredients [476, 526]. Inhomogeneity [451]. Initial [189]. Injury [320]. Installation [257]. Insurance [432, 437, 304]. Integer [85]. Integral [461]. Integrated [599, 480, 456]. Integration [543, 571]. Intensities [457]. Intensity [109, 115]. Inter [581]. Inter-NFHS [581]. Interaction [539, 242]. Interactions [8, 591]. Interactive [496]. Interdistributional [572]. Interest [74, 80]. Internal [379]. International [443, 100, 112, 120, 138, 152, 170, 180, 191, 200, 211, 228, 235, 247, 262, 275, 293, 301, 314, 331, 341, 351, 361, 373, 384, 399, 412, 427, 436, 445, 454, 464, 472, 478, 486, 493, 503, 514, 522, 534, 545, 560, 569, 576, 582, 588, 594, 600, 606, 613, 619]. Internet [365]. Interpolated [521]. Interpolation [433]. Interruptions [86]. Intersections [378]. Interval [267, 344, 565, 521, 505, 400, 321, 62, 334, 473, 104]. Interval-Valued [473]. Intervals [35, 391, 579, 465, 47]. Intervention [442]. Intra [105]. Intra-Veinous [105]. Introduced [335, 355, 405, 439]. Introduction [508]. Introductory [362]. Invariance [268, 234]. Invariant [414]. Inverse [423, 82, 578, 322, 205, 255]. Inverted [495, 511, 611]. Investigation [19]. Investigations [390]. Investment [531, 137]. Invitation [161]. Involving [12]. Iraq [5]. Issues [490]. Items [28]. Iterated [485]. Iterations [549]. Iterative [424]. IV [355].

Jaundice [125]. Join [123]. Joint [520, 416, 124]. Journal [100, 112, 120, 138, 152, 170, 180, 191, 200, 211, 228, 235, 247, 262, 275, 293, 301, 314, 331, 341, 351, 361, 373, 384, 399, 412, 427, 436, 445, 454, 464, 472, 478, 486, 493, 503, 514, 522, 534, 545, 560, 569, 576, 582, 588, 594, 600, 606, 613, 619]. Journals [122]. Jump [300]. Justification [263].

Kelantan [602]. Kendal [203]. Kendall [359]. Kenya [567, 444]. Kernel [194, 84, 187]. Kidnappings [528]. Knot [107]. Knots [107]. Knowledge [265]. Known [299, 424, 101]. Kolmogorov [40, 310]. Kronecker [476, 526]. Kumaraswamy [106, 281, 330, 160, 198]. Kumaraswamy-Half-Cauchy [160]. Kurtosis [220, 67].

Lack [450]. Lack-of-fit [450]. Lambda [48]. Land [500]. Lanka [365]. Laplace [10]. Large [518, 529, 437, 97, 128, 130, 210, 169]. Large-Scale [169]. Lasso [151, 240]. Last [289]. Last-order [289]. Latent [523, 102, 287, 377, 69, 268, 539, 318, 135, 242]. Law [210]. Lawley [395]. Lead [221]. Learning [468, 353, 249, 616, 414, 387, 531]. Least [268, 403]. Left [11]. Legal [438]. Lehmann [585, 536]. Lemma [563]. Length [46]. Leone [488, 260, 615, 611]. Leone-G [615]. Lev [467]. Level [122, 591].

Levels [567]. Leverage [614]. Libby [95]. Life [43, 536, 315, 334]. Lifestyle [350]. Lifetime [54, 141, 368, 181, 584, 598, 294, 332, 132, 176, 178, 578]. Lifetimes [421]. Likelihood [555, 213, 415, 44, 552, 20, 535, 129, 155, 104, 410]. Likelihood-based [20]. Likelihoods [248]. Limit [72, 251, 490, 250]. Limitations [491]. Lindley [368, 455, 337, 259, 482, 55, 418, 550, 145]. Lindley- [455]. Lindley-Burr [550]. Lindley-type [337]. Linear [555, 157, 239, 328, 474, 287, 182, 151, 241, 118, 153, 460, 484, 73, 408, 597, 133, 308, 117, 282, 71, 75, 30, 184, 291, 32, 129, 410, 356]. Linearly [66]. LINEX [477]. Linkage [27]. Links [463]. Linnik [398]. Lipschitz [540]. Litterman [523]. Liu [460, 1]. Load [489]. loaded [264]. Loadings [173]. Local [372, 90, 422]. Locality [130]. Location [88]. Log [164, 146, 417, 404, 181, 45, 389, 541, 294, 357, 364, 306]. Log-Density [306]. Log-Logistic [404, 389, 541, 294, 357]. Log-McBurr [181]. Log-Normal [164, 146, 45]. Logarithmic [380]. Logical [219]. Logistic [417, 396, 404, 336, 204, 499, 285, 601, 389, 532, 541, 326, 278, 294, 357]. Logistic-G [532]. Logit [567]. Lomax [383, 459, 469, 553]. Long [555, 204]. Long-term [204]. Longitudinal [162, 153, 416, 393, 19]. Lord [362]. Loss [10, 324, 55, 477, 140]. Low [411]. Lower [374]. LSEs [512]. Lyapunov [236].

M [340, 413, 470, 340, 413, 470]. M/M/1 [340, 413, 470]. Machine [387]. Macro [172]. Macroeconomic [513]. Main [86]. Maintenance [165]. Making [208]. Malnutrition [5]. Management [379, 602]. Managerial [168]. Mann [589, 291]. MANOVA [395]. Maps [35]. Margin [292]. Marginal [451, 162, 381, 20, 364, 108, 49]. Marginalized [73, 552]. Marginals [256]. Margins [272]. Mark [300]. Market [468, 328, 243, 29, 2]. Markets [279]. Markov [448, 328, 377, 110, 579, 186, 8, 430, 442, 548, 480, 80]. Marks [458]. Marshall [488, 179, 89, 45, 469, 256, 532, 570]. Marshll [322]. Massive [306]. Matched [381]. Matched-Pair [381]. Materials [525]. Mathematics [345]. Matrices [25]. Matrix [296, 512, 26]. Max [33]. Max-stable [33]. Maximal [91, 115]. Maximization [533]. Maximizing [61]. Maximum [555, 237, 72, 44, 552, 193, 104]. Maxwell [504, 419, 543, 549]. Mc [6]. Mc- [6]. McBurr [181]. McDonald [53]. MCMC [15, 352]. McNemar [255]. MCP [520]. MCP-Mod [520]. Mean [607, 344, 299, 56, 18, 8, 223, 509, 134, 540, 113, 465]. Mean-correction [223]. Mean-Field [540, 8]. Mean-Variance [56]. Meanings [39]. Means [164, 146, 491, 4]. Measure [128, 130, 90, 461, 422, 7, 586, 49]. Measurement [441, 320, 31, 288, 28, 356]. Measurements [409]. Measures [572, 398, 566, 115, 96, 145, 234, 4]. Measuring [451, 77, 51]. Mechanical [419, 543, 549]. Median [22, 521]. Medical [380, 442]. Mellitus [483, 342]. Memory [555, 109]. Merton [300]. Meta [587, 593, 618]. Meta-Analysis [618, 587, 593]. Method [401, 575, 504, 98, 230, 415, 394, 254, 345, 111, 81, 117, 326, 124, 193, 199, 402].

Methodological [16]. Methodologies [420]. Methodology [544, 214, 466, 306]. Methods [183, 352, 432, 270, 601, 309, 107, 551, 108, 617]. Mexican [483]. Mexico [499, 528]. MFPT [524]. Microfinance [379]. Minimal [147]. Minimax [155]. Minimizing [208]. Minimum [147, 447, 589]. Mining [111, 363]. miRNAs [136]. mIRT [288]. Misclassified [371]. Mises [188]. Missing [509]. Mitigating [264]. Mixed [151, 603, 254, 73, 75, 305, 199]. Mixture [15, 21, 82, 153, 476, 526, 274, 53, 546, 547, 615, 140]. Mixtures [317, 524]. Mod [520]. Mode [12]. Model [106, 74, 10, 468, 368, 404, 538, 260, 474, 483, 203, 287, 491, 139, 109, 350, 411, 370, 499, 300, 285, 273, 342, 55, 246, 175, 151, 127, 216, 359, 263, 186, 19, 372, 91, 476, 526, 158, 266, 616, 425, 557, 389, 343, 550, 1, 105, 430, 388, 484, 567, 131, 223, 546, 547, 548, 133, 145, 456, 340, 413, 313, 539, 296, 117, 364, 282, 28, 173, 326, 440, 463, 48, 7, 269, 551, 403, 473, 283, 101, 140, 215, 129, 104, 41, 375]. Model-based [313]. Modeling [380, 325, 239, 147, 324, 607, 109, 3, 599, 365, 562, 153, 279, 510, 418, 268, 266, 345, 574, 392, 124, 489, 242, 2]. Modelling [533, 528, 69, 241, 520, 481, 475, 85, 525, 354, 46]. Models [252, 328, 353, 336, 182, 213, 162, 181, 338, 237, 70, 202, 590, 605, 599, 556, 603, 249, 256, 448, 311, 274, 13, 78, 460, 197, 530, 125, 99, 73, 416, 220, 471, 597, 450, 38, 382, 117, 604, 288, 390, 75, 305, 516, 184, 535, 137, 410, 356]. Modified [22, 575, 407, 248, 246, 118, 357, 382, 148]. Modulated [271]. Modulo [339]. Molecular [425]. Moment [394, 428]. Moment-Based [428]. Moments [572, 374, 223]. Money [295]. Monitor [19]. Monitoring [554, 261]. Monotone [156, 509]. Monotonic [304]. Monotonicity [303]. Monoxide [392]. Monte [432, 67]. Monthly [259]. Moran [349]. Morse [461]. Morse-Transue [461]. Mortality [350, 304, 480]. Motion [225, 32]. Motor [437]. movements [443]. Moving [599, 480, 456]. MQCIA [148]. Much [581]. Multi [416, 68, 234, 4, 168, 375]. Multi-Dimensional [68, 234]. Multi-Factor [375]. Multi-stages [168]. Multi-state [416]. Multi-way [4]. Multicollinearity [554]. Multidimensional [574]. Multilevel [237, 69, 73]. Multimorbidity [387]. Multinomial [23, 381]. Multiperiod [56]. Multiple [353, 27, 566, 83, 309, 347, 43, 31, 61, 516]. Multiple-Objective [31]. Multiply [164]. Multivariable [440]. Multivariate [523, 316, 398, 415, 347, 44, 90, 276, 303, 220, 47, 133, 92, 108]. Myeloma [43].

N [214]. N-Point [214]. Nadarajah [449]. Nadaraya [134]. Naive [468]. Namibia [574]. National [27, 161]. Natural [467]. Nearest [485]. Neck [495]. Negative [149, 425, 132, 612]. Negatively [210, 251]. Negatives [140]. Neighbourhood [98]. Nelson [277]. Nephropathy [11, 59]. Nested [53]. Network [151, 599, 502, 319]. Networks [123, 97, 527, 88, 196]. Neural [599, 88, 196]. Newborns [125]. NFHS [581]. NHPP [404]. Nigeria [437, 544, 599, 392, 480, 463, 297]. Nine [617]. No [100, 112, 120, 138, 152, 170, 180, 191, 200, 211, 228, 235, 247, 262, 275, 293, 301, 314, 331, 341, 351, 361, 373, 384, 399, 412, 427, 436, 445, 454, 464, 472, 478, 486, 493, 503, 514, 522, 534, 545, 560, 569, 576, 582, 588, 594, 600, 606, 613, 619]. Node [345]. Noisy [303]. Non [396, 328, 352, 434, 300, 311, 527, 382, 296, 540]. Non-graded [527]. Non-Homogeneous [396]. Non-Identifiability [296]. Non-Lipschitz [540]. Non-Parametric [434]. Non-standard [352, 382]. Non-Sum [311]. Nonchronic [556]. Noncompliance [81]. Nonlinear [182, 603, 8, 616, 116]. Nonnegative [512]. Nonnegative-Definite [512]. Nonparallel [143]. Nonparametric [9, 521, 72, 87, 303, 530, 440, 104]. Nonprobability [571]. Nonrandomized [309]. Nonstationary [44]. Normal [164, 146, 592, 344, 299, 126, 471, 433, 63, 45]. Normality [116]. Note [54, 25, 166, 182, 60, 16]. Notion [346]. Novel [167]. Novick [95]. Nuisance [20]. Number [174, 516]. Numbers [210]. Numerical [549, 420].

O [461]. O-dot [461]. Objective [273, 31]. Observations [84, 303]. Observers [515]. Obstetric [338]. Obtained [504, 452]. Occurrence [135]. Odd [585, 368, 455, 418]. Oil [591]. Older [115]. Olkin [179, 488, 89, 322, 45, 469, 256, 532, 570]. Olkin- [488]. Omission [597]. One [219, 395, 105, 509, 20]. One-sided [20]. One-Way [395]. Ones [66]. Onset [59]. Open [283]. Operating [524]. Operational [506]. Optimal [367, 119, 171, 476, 526, 438, 577, 47, 474]. Optimization [252, 103, 564]. Optimized [376]. Optimum [206]. Options [447]. Order [15, 21, 163, 182, 126, 214, 223, 20, 63, 269, 18, 289]. Ordered [216, 66, 28]. Orders [618]. Ordinal [49]. Ordinary [403]. ordination [39]. Ornstein [250]. Other [115]. Outlier [317]. Outliers [401, 230, 71]. Ovarian [240]. Over-dispersed [274]. Overall [240]. Overcoming [491]. Overdispersion [53]. Overlapping [66]. Oxygen [115].

P [608]. P-values [608]. Pair [381]. Paired [7]. Panel [204, 528]. Paradox [479, 362]. Parameter [10, 368, 239, 252, 230, 513, 394, 249, 507, 308, 552, 173, 580, 205, 142, 176]. Parameters [494, 421, 103, 477, 321, 561, 552, 50, 62, 20, 63, 580, 233]. Parametric [336, 287, 434, 338, 33, 11, 400, 530, 420]. Parasites [475]. Pareto [226, 407, 322, 205, 14, 286]. Parliamentarians [481]. Parsimonious [359, 605]. Partial [268, 410, 280]. Partially [171, 460, 424, 144, 184, 101, 356]. Particle [252, 249]. Partitioning [457, 388]. Parts [66]. Passage [458, 253]. Path [268]. Pathogenicity [218]. Pathwise [105]. Patients [338, 59, 31, 240, 305]. Patrol [257]. Pattern [414]. Patterns [5, 349]. Payable [199]. PBDE [415]. Peak [115]. Pearson [359]. Pedestrian [369]. Penalized [64, 519, 107]. Penetrans [567]. Pensions [77]. Pentuals [508]. Percentages [492]. Performance [533, 391, 304, 81]. Period [507, 581, 580, 305]. Periodic [165]. Perspective [533, 284]. Perspectives [39]. Petroleum [245]. Pharmaceutical [137]. Phase [46, 124]. Phase-type [46]. Phototherapy [125]. Physiologic [92]. Piecewise [225]. Places [88]. Plans [334]. Platform [257, 496]. Plausible [607]. Pleiotropic [96]. plot [577]. Plots [597, 144]. Point [566, 214, 505, 349, 193, 233]. Points [83, 103, 189, 537, 614]. Poisson [423, 417, 396, 267, 102, 504, 190, 139, 578, 584, 482, 55, 175, 394, 449, 385, 366, 13, 558, 1, 598, 536, 85, 463, 305, 201, 178]. Poisson-G [584]. Poisson-Log-Logistic [417]. Policy [358]. Polya [524]. Polynomial [372, 339, 450, 433]. Polynomially [121]. Polynomials [119]. Population [483, 434, 304, 358]. Populations [299, 521, 246]. Portion [199]. Positioning [345]. Positivity [303]. Possibility [52]. Possibly [304]. Post [320, 233]. Post-change [233]. Post-Hospital [320]. Posterior [333]. Poverty [34, 574]. Power [585, 467, 374, 578, 584, 259, 45, 482, 469, 418, 329, 61, 511, 553, 591, 67, 402]. Pradesh [581, 492]. Precise [127]. Predict [483, 499]. Predicting [240, 444, 193]. Prediction [15, 575, 607, 480, 47, 144]. Predictive [151, 232, 617]. Predictor [597]. Preference [463]. Preferences [353, 431]. Preferential [97, 128]. Premium [55]. Premiums [80]. Prescription [115]. Presence [290, 441, 72, 20, 282, 196]. Preserving [607, 129]. Pressure [441, 425]. Previsions [431]. Price [77, 386]. Pricing [437]. Primary [608]. Prime [397, 91]. Primes [339, 91]. Principal [367, 81, 426, 466, 184]. Principle [128, 221]. Prior [232, 75]. Probabilistic [91, 41]. Probabilities [462, 311]. Probability [592, 156, 219, 84, 174, 218, 225, 583, 559, 418, 311, 571, 80, 93, 79, 200, 211, 301, 314, 331, 341, 100, 112, 120, 138, 152, 170, 180, 191, 228, 235, 247, 262, 275, 293, 351, 361, 373, 384, 399, 412, 427, 436, 445, 454, 464, 472, 478, 486, 493, 503, 514, 522, 534, 545, 560, 569, 576, 582, 588, 594, 600, 606, 613, 619]. Problem [245, 485, 40, 429, 531]. Problems [25, 377, 529, 559, 457, 376, 118]. Procedure [521, 18]. Proceedings [438]. Process [396, 23, 300, 44, 37, 466, 554, 80, 250]. Processes [555, 253, 239, 316, 166, 33, 238, 68]. Product [461]. Production [386]. Products [114]. Professional [12]. Profile [248]. Programming [118]. Programs [227, 276, 264]. Progression [305]. Progressive [494, 421, 334]. Project [168, 137]. Projection [358]. Proof [419, 543, 549]. Propensity [544]. Properties [179, 54, 141, 368, 154, 459, 538, 407, 190, 270, 435, 578, 469, 207, 418, 389, 343, 550, 6, 406, 598, 511, 553, 536, 185, 393]. Proportional [4]. Proportionate [229]. Proportions [9, 304]. Proposed [347]. Proposition [30]. Prostate [193]. Proxies [115]. Proxy [115]. Psychological [219].

QCIA [148]. Qu [76]. Quadratic [408, 24]. Quality [27, 441, 587]. Quality-Asthma [587]. Quantification [515]. Quantile [64, 575, 76, 87, 158, 144, 586, 297]. Quantiles [565, 187]. Quarantine [618]. Quartile [230]. Quasi [44]. Question [452]. Questioning [309]. Queue [319]. Queueing [86]. Quick [118, 148, 189, 148]. Quinn [30].

R [539, 310]. R&D [168, 137]. R-Software [310]. Rainfall [259, 599, 444].

Random [346, 42, 98, 97, 128, 130, 421, 110, 210, 251, 202, 366, 78, 577, 461, 431, 16, 93, 79, 334, 137, 272]. Randomized [485, 81]. Range [12, 159]. Rank [407, 269]. Ranking [265, 515]. Rasch [320]. Rate [74, 194, 538, 77, 213, 83, 72, 590, 505, 408, 142, 525, 264]. Rates [17, 304, 480]. Ratio [265, 213, 484, 535, 167, 193, 129, 155]. Rationality [447]. Raw [525]. Rayleigh [21]. Real [330, 496]. Reasoning [52, 62, 169]. Receiver [524]. Recently [335, 355, 405, 439]. Recognition [177, 414]. Record [374, 126]. Recovery [338]. Recurrence [374]. Recursive [186, 197]. Reduction [4]. Reference [483]. Referring [346]. Refines [376]. Reflecting [32]. Regimes [328]. Regions [47]. Regression [555, 573, 157, 64, 194, 260, 353, 287, 139, 434, 411, 204, 285, 601, 342, 202, 175, 347, 556, 119, 372, 418, 87, 158, 616, 519, 530, 1, 107, 47, 133, 450, 38, 134, 591, 71, 440, 463, 30, 551, 516, 184, 473, 535, 614, 410, 297, 280]. Regressions [76, 143]. Regressive [480, 448]. Regularized [158]. Regulatory [94]. Rehabilitation [320]. Reinsurance [437]. Reinsurer [199]. Related [397, 110, 160]. Relations [374]. Relationship [102, 395]. Relationships [94, 92]. Relative [150]. Reliability [380, 212, 404, 244, 370, 409, 217, 6, 321, 145, 271, 587, 593]. Reliable [302]. Religion [463]. Removals [421, 334]. Repeated [409, 31]. Replacement [79]. Replicated [171, 144]. Replications [450]. Reproducibility [618]. Reproducible [467]. Reproductive [595]. Research [17, 363, 51]. Reset [447]. Residual [533, 141, 597]. Residuals [555]. Resolution [222]. Resolving [25]. Resource [602]. Response [309, 214, 31, 466, 28]. Response-Adaptive [31]. Ressel [203]. Restricted [347, 551, 356]. Restrictions [117]. Result [130]. Results [467, 337, 160]. Retail [88]. Return [77]. Returns [266]. Reversed [141]. Review [609, 601, 420, 284]. Reviewer [100, 112, 120, 138, 152, 170, 180, 191, 200, 211, 228, 235, 247, 262, 275, 293, 301, 314, 331, 341, 351, 361, 373, 384, 399, 412, 427, 436, 445, 454, 464, 472, 478, 486, 493, 503, 514, 522, 534, 545, 560, 569, 576, 582, 588, 594, 600, 606, 613, 619]. Revisit [501]. Reweighted [134]. Right [72, 11]. Risk [483, 208, 432, 42, 379, 290, 246, 18, 243, 378, 80]. Risks [277]. River [602]. Rivers [297]. Road [246, 224, 283]. Roadways [411]. Robust [22, 285, 517, 136, 26]. ROC [248]. Role [468]. Roles [463]. Root [238]. Roots [327, 339]. Rough [122]. RSP [150]. Ruin [80]. Rules [139, 189]. Ruminal [603]. Runs [171, 66]. Rural [379, 378].

Saddlepoint [121]. Safe [56]. Safety [520, 506, 124]. Sample [529, 438, 507]. Samplers [185]. Samples [469, 457, 571, 50]. Sampling [106, 33, 206, 79, 525, 610, 334, 255, 172]. SARIMA [599]. SARS [610]. SARS-CoV-2 [610]. SAS [172]. Scale [10, 236, 169, 135, 167]. Scaled [318]. Scales [209]. Scans [185]. Scenario [551]. Scheme [494, 421, 300, 308]. School [353]. Sci [122]. Sci-tech [122]. Score [244, 544, 13]. Scoring [150]. SDE [113]. Seasonal [599, 241]. Second [182, 18, 214, 476, 526].

Second-Order [214, 18]. Sectional [528]. Sections [506]. Security [429]. Seeding [161]. Seemingly [535]. Selected [526]. Selecting [285, 88]. Selection [139, 616, 388, 107, 375]. Self [109, 333, 285]. Self-Exciting [109]. Self-information [333]. Self-Selecting [285]. Semi [328, 530]. Semiparametric [336, 162, 70, 13, 460, 184, 167, 193]. Sequences [397, 16]. Sequential [508, 462, 249, 409, 205, 233]. Series [114, 325, 401, 500, 566, 528, 236, 609, 327, 241, 263, 153, 197, 456]. Server [86, 217, 271]. Service [123, 257, 340, 413, 470]. Set [210, 251, 122, 117]. Set-to-Zero [117]. Set-Valued [210, 251]. Sets [564]. Several [70]. Severity [324, 369, 378, 567]. Sex [463]. Shape [507, 205, 551]. Shaped [538, 142, 176]. Shared [252, 338]. Shift [586]. Shifted [226, 457]. Shifted-Geometric [457]. Ship [41]. Ship-Bridge [41]. Short [25]. Shortest [344]. Showing [234]. Shuttling [50]. Sibuya [279]. Sibuya-type [279]. sided [20]. Sidewalks [369]. Sidhu [353]. Sigma [93]. Sigma-Complete [93]. Sign [376]. Signalized [378]. Significance [182, 258, 218, 231]. Simple [22, 106, 491, 329, 185, 610]. Simulation [15, 123, 238, 495, 564, 425, 53, 107]. Simulation-Based [564]. Simultaneous [381, 35, 465, 296, 440]. Single [64, 158]. Single-Index [64]. Singly [146]. Singular [426]. Singularity [25, 296]. Sinh [495]. Sinister [199]. SIS [548]. Site [575]. Situ [602]. Size [374, 438, 586]. Size-Biased [374]. Skew [126, 471]. Skew-Normal [126, 471]. Skewed [497]. Slope [476, 526]. Small [35, 507]. Smeed [246]. Smirnov [40]. Smooth [227, 276, 498]. Smoothing [60]. Snalysis [273]. Socioeconomic [69]. Software [111, 310]. Solution [218, 127, 113]. Solutions [377]. Solve [485]. Solving [118]. Some [312, 141, 154, 65, 203, 377, 435, 601, 110, 337, 274, 385, 185, 516]. Space [249, 8]. Spaced [107]. Spacing [193]. Sparse [240, 329, 452]. Spatial [33, 5, 19, 296, 172]. Spatial-longitudinal [19]. Spatio [349]. Spatio-temporal [349]. SPBVAR [443]. Spearman [359, 90]. Spearman-Type [90]. Species [102]. Species-Area [102]. Speed [151, 419, 543, 549]. Spheres [50]. Spherical [214]. Spline [107, 440, 410]. Split [577]. Split-plot [577]. SPRE [167, 193]. Square [451, 82, 216, 605, 364, 422, 49, 215, 113]. squared [539]. Squares [268, 117, 403]. Sri [365]. Stabilization [32]. Stable [263, 33]. Stage [441, 69, 18]. Stages [442, 168]. Staggered [505]. Stand [86]. Stand-By [86]. standard [352, 382]. Standardization [601]. Standardized [278]. Star [171]. State [239, 249, 8, 448, 297, 416]. State-Space [249]. States [556, 516]. Statistic [395, 465, 349, 188]. Statistical [106, 538, 177, 203, 259, 34, 490, 502, 44, 19, 481, 362, 339, 510, 386, 6, 96, 329, 554, 67, 593, 261, 231]. Statistics [15, 21, 163, 126, 12, 289, 509, 63, 100, 112, 120, 138, 152, 170, 180, 191, 200, 211, 228, 235, 247, 262, 275, 293, 301, 314, 331, 341, 351, 361, 373, 384, 399, 412, 427, 436, 445, 454, 464, 472, 478, 486, 493, 503, 514, 522, 534, 545, 560, 569, 576, 582, 588, 594, 600, 606, 613, 619]. Status [565, 69]. Stay [618, 46]. Stay-at-Home [618]. Steel [466, 554]. Step [115, 315, 509].

Step-Stress [315]. Stochastic [523, 77, 33, 3, 557, 223, 308, 58, 489, 32, 356, 2]. Stock [468, 266, 375]. Stop [378]. Stop-Controlled [378]. Stores [88]. Strategies [56, 295]. Stratification [81]. Stratified [568]. Strength [370, 321]. Stress [106, 370, 321, 315]. Stress-Strength [370, 321]. Strings [66]. Strong [210]. Strongly [68]. Structural [533, 607, 3, 76, 539, 390, 242]. Structure [491, 55, 243, 268, 229, 57, 234, 75, 512, 166]. Structured [591]. Structures [474, 129]. STSP [55]. Student [39, 169, 46]. Students [353, 556]. Studies [603, 409, 124, 587, 618, 41]. Study [183, 212, 396, 379, 259, 415, 148, 107, 57, 67, 402, 365]. Style [353]. Sub [246, 465, 577, 403]. Sub-mean [465]. Sub-Model [403]. Sub-plot [577]. Sub-Populations [246]. Subclassification [517]. Subject [409]. Subjective [219]. Subsidy [264]. Such [318]. Sum [311, 79, 117]. Sum-to-Zero [117]. Summarizing [363]. Sums [110, 366]. Supply [245]. Support [237]. Surface [500, 214, 524, 466]. Surprising [185]. Survey [595, 541, 571, 617]. Survival [538, 11, 45, 225, 69, 323, 415, 453, 240, 420, 416, 193, 286]. Susceptibility [14]. Sustained [12]. SVM [375]. Swap [245]. Swarm [252]. Switch [56]. Switch-When-Safe [56]. Symbolic [473]. Symmetric [559]. Symmetrical [521]. Symmetry [216, 605, 422, 215]. Synchronization [443]. System [165, 441]. Systems [86, 32].

t [497]. t-test [497]. Table [272]. Tables [451, 216, 359, 605, 169, 364, 422, 4, 49, 215]. Tail [563]. tailed [324, 428]. Taking [483]. Task [271]. Tau [229, 28]. Tau-equivalent [28]. Tau-Proportionate [229]. Taxi [264]. TB [305]. Teaching [12]. tech [122]. Techniques [267, 177, 323, 175, 428]. Technology [288]. Technology-Enhanced [288]. Telecommunication [365]. Temperature [147, 500]. Tempering [466]. Temporal [183, 349]. term [204]. Terms [333]. Terry [7]. Test [287, 40, 258, 395, 421, 76, 521, 347, 376, 465, 268, 115, 546, 547, 589, 509, 20, 234, 67, 334, 255, 535, 233, 129, 497]. Testing [182, 381, 218, 347, 409, 87, 99, 116, 61, 133, 450, 20, 315, 440]. Tests [9, 213, 13, 303, 155, 272]. Text [363]. Texture [515]. Texture-Ranking [515]. Textures [515]. Their [214, 475, 337]. Theorem [97, 72, 263, 403, 94]. Theorems [251]. Theoretical [423]. Theory [585, 52, 316, 362, 254, 37, 406, 442, 444, 199]. There [586]. Three [244, 69, 476, 111, 507, 552, 580, 242, 402]. Three-arm [402]. Three-Parameter [507, 580]. Three-way [242]. Threshold [57]. Threshold-Free [57]. Thresholds [537]. Time [114, 325, 253, 401, 239, 500, 377, 209, 566, 338, 33, 421, 528, 236, 609, 11, 59, 225, 327, 263, 153, 31, 197, 496, 430, 99, 536, 456, 308, 568, 166]. Time-Series [153]. Time-Stratified [568]. Time-to-Event [239]. Time-to-Recovery [338]. Time-Varying [253, 166]. Times [108]. Tool [220]. Tools [566]. Topp [260, 488, 615, 611]. TOPSIS [122]. Total [434].

Tough [467]. Tournament [161, 161]. Trace [395]. Trading [295]. Traditional [376]. Traffic [411, 246, 224, 257, 506, 283]. Trajectories [153]. Transactions [502]. Transcriptional [94]. Transfer [414, 456]. Transform [127, 364]. Transformation [267]. Transformed [82, 497]. Transformed/Inverse [82]. Translationally [414]. Transmetric [207, 195]. Transmission [430]. Transmuted [179, 407, 259, 342, 198]. Transue [461]. Treatment [367, 391, 81, 305]. Treatment/Condition [367]. Tree [111, 517]. Trees [524, 414]. Trend [9, 241, 365, 568]. Trends [259, 363]. Trial [81]. Trials [292, 17, 261]. Triangle [140]. Truels [462]. Truncated [102, 336, 11, 482, 302, 598, 332]. Tunga [567]. Turkish [204]. Tutorial [539]. Tweedie [467]. Two [164, 146, 15, 21, 368, 230, 439, 359, 18, 60, 465, 319, 81, 509, 39, 142, 176, 358, 512]. Two-child [358]. Two-Dimensional [60]. Two-Parameter [368, 230, 142, 176]. Two-stage [18]. Two-step [509]. Two-Way [359]. Type [89, 585, 483, 265, 109, 59, 605, 465, 90, 561, 536, 611, 174, 337, 279, 46, 494, 396, 336, 421, 342, 506, 334]. Type-A [506]. Type-I [561, 396]. Type-II [561, 611, 494]. Typed [130]. Types [98].

Uhlenbeck [250]. UK [27]. Ulcers [441]. Unbiased [157, 115]. Unbiasedness [263]. Uncertain [218]. Uncertainty [52, 350, 137]. Under-5 [480]. Underdispersed [558]. Underlying [135]. Underreporting [290]. Understand [19]. Unequal [457]. Ungauged [575]. Unified [572, 559]. Uniform [462, 37, 570]. Uniqueness [58]. Unit [238, 327]. United [556, 516]. University [46, 353]. Unknown [321]. Unrelated [452, 535]. Unreliable [340, 413, 470]. Unsupervised [387]. Urban [378]. Use [189, 317, 206, 492, 551, 297]. Used [602, 420]. Using [592, 10, 267, 102, 352, 265, 109, 544, 204, 441, 521, 259, 513, 246, 590, 359, 477, 524, 150, 448, 43, 266, 349, 574, 506, 442, 329, 159, 597, 456, 554, 539, 63, 604, 88, 390, 85, 463, 167, 193, 46, 261, 323, 279]. Utilisation [354]. Utilization [525]. Uttar [581, 492].

Vacation [413, 470]. Validity [409]. Value [147, 42, 437, 256, 254, 589, 444, 426, 199]. Valued [210, 251, 473]. Values [374, 607, 126, 608]. Variable [102, 177, 265, 447, 616, 135, 150]. Variables [183, 42, 377, 143, 210, 251, 513, 202, 206, 268, 366, 527, 115, 92, 16, 93, 79, 318]. Variance [141, 467, 367, 299, 56, 360, 144]. Variances [258, 424, 318]. Variant [218, 155]. Variation [35, 63, 135]. Various [59, 185]. Varying [253, 377, 460, 308, 166]. Vascular [27]. Vasicek [74]. VECM [513]. Vector [237, 263, 591]. Vectors [465, 509, 24]. Vehicle [369]. Vehicle-Pedestrian [369]. Veinous [105]. Version [383, 28]. Versus [530]. Via [15, 126, 163, 139, 236, 616, 227, 276, 388, 517, 571]. Views [280]. VIF [150]. Vine [408, 596]. Violence [544]. Viral [489]. Virus [510]. Visual [566]. Visualizing [448, 496, 133]. Visually [451]. Vol [100, 112, 120, 138, 152, 170, 180, 191, 200, 211, 228, 235, 247, 262, 275, 293, 301, 314, 331, 341, 351, 361, 373, 384, 399, 412, 427, 436, 445, 454, 464, 472, 478, 486, 493, 503, 514, 522, 534, 545, 560, 569, 576, 582, 588, 594, 600, 606, 613, 619]. Volatilities [2]. Volatility [523, 3, 223]. Volume [114, 411]. Volumes [411]. Vulnerable [297].

Wald [67]. Waste [485]. Water [602, 425, 591]. Watson [134]. Way [395, 359, 318, 242, 4]. Weak [42]. Wear [507, 580]. Wear-Out [507, 580]. Weaving [506]. Web [496]. Web-Based [496]. Weekly [568]. Weibull [15, 149, 106, 212, 423, 368, 494, 154, 226, 260, 330, 370, 421, 578, 342, 307, 477, 389, 343, 406, 561, 302, 198, 598, 511, 507, 357, 552, 282, 580]. Weibull- [343, 406, 307, 302]. Weighted [212, 572, 558, 480, 145, 284, 28, 393]. Weights [571, 617]. Well [431]. Well-Behaved [431]. Whilst [69]. Whitney [589, 291]. Whittle [555]. Whole [577]. Wilcoxon [589, 291]. Wilkes [353]. Winning [462]. Wire [466]. Wishart [501]. Within [492]. Without [450, 79]. WLSEs [512]. Wold [263]. Wolfowitz [67]. Women [595, 481, 492, 463]. Wood [114]. Working [413, 470]. Wrapped [190, 201]. WSN [345].

X [149]. XII [178, 149, 573, 89, 181, 550, 132].

Years [69].

Zero [102, 190, 609, 482, 175, 13, 598, 117, 463, 201]. Zero-Inflated [609, 463, 190, 175, 13]. Zeros [140, 354]. Zone [246]. Zurbenko [310].

References

Mansson:2012:ILE

[1] Kristofer Mansson, B. M. Golam Kibria, Par Sjolander, and Ghazi Shukur. Improved Liu estimators for the Poisson regression model. International Journal of Statistics and Probability, 1(1):2–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

     ccsenet.org/journal/index.php/ijsp/article/view/16669.

Zheng:2012:SME

[2] Jin Zheng and Dejun Xie. Stochastic modeling and estimation of market volatilities with applications in financial forecasting. International Journal of Statistics and Probability, 1(1):7–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

     org/journal/index.php/ijsp/article/view/14970.

Escobar:2012:SCM

[3] Marcos Escobar, Tim Friederich, Mikhail Krayzler, Luis Seco, and Rudi Zagst. Structural credit modeling under stochastic volatility. International Journal of Statistics and Probability, 1(1):20–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/16674.

Yamamoto:2012:MPR

[4] Kouji Yamamoto, Yuri Nozaki, and Sadao Tomizawa. Measures on proportional reduction in error by arithmetic, geometric and harmonic means for multi-way contingency tables. International Journal of Statistics and Probability, 1(1):36–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/13869.

Khamis:2012:ABS

[5] Faisal G. Khamis and Ghaleb A. El-Refae. Association between spatial patterns of acute malnutrition and household income in Iraq — 2004. International Journal of Statistics and Probability, 1(1):43–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/14848.

Marciano:2012:MDS

[6] Francisco W. P. Marciano, A. D. C. Nascimento, M. Santos-Neto, and G. M. Cordeiro. The Mc-Γ distribution and its statistical properties: An application to reliability data. International Journal of Statistics and Probability, 1(1):53–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/16678.

Tahata:2012:MDE

[7] Kouji Tahata, Keigo Kozai, and Sadao Tomizawa. Measure of departure from extended Bradley–Terry model for paired comparisons. International Journal of Statistics and Probability, 1(1):72–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/16681.

Kolokoltsov:2012:NMG

[8] Vassili N. Kolokoltsov. Nonlinear Markov games on a finite state space (mean-field and binary interactions). International Journal of Statistics and Probability, 1(1):77–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/16682.

Alvo:2012:NTT

[9] Mayer Alvo and Marie-Pascal Berthelot. Nonparametric tests of trend for proportions. International Journal of Statistics and Probability, 1 (1):92–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/16683.

Ali:2012:SPE

[10] Sajid Ali, Muhammad Aslam, Nasir Abbas, and Syed Mohsin Ali Kazmi. Scale parameter estimation of the Laplace model using different asymmetric loss functions. International Journal of Statistics and Probability, 1(1):105–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/14429.

Grover:2012:PAE

[11] Gurprit Grover and Alka Sabharwal. A parametric approach to estimate survival time of diabetic nephropathy with left truncated and right censored data. International Journal of Statistics and Probability, 1(1): 128–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/14522.

Jacobbe:2012:ISP

[12] Tim Jacobbe and Bob Horton. The importance of sustained professional development for teaching statistics — an example involving the mode and range. International Journal of Statistics and Probability, 1 (1):138–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/16684.

Li:2012:STS

[13] Chin-Shang Li. Score tests for semiparametric zero-inflated Poisson models. International Journal of Statistics and Probability, 1(2):1– ??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/16099.

Zea:2012:BEP

[14] Luz Milena Zea, Rodrigo B. Silva, Marcelo Bourguignon, Andrea M. Santos, and Gauss M. Cordeiro. The beta exponentiated Pareto distribution with application to bladder cancer susceptibility. International Journal of Statistics and Probability, 1(2):8–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/16301.

Abushal:2012:BPB

[15] Tahani Abdulrahmaan Abushal and Areej M. Al-Zaydi. Bayesian prediction based on generalized order statistics from a mixture of two exponentiated Weibull distribution via MCMC simulation. International Journal of Statistics and Probability, 1(2):20–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/16092.

Rambaud:2012:MNC

[16] Salvador Cruz Rambaud and Antonio Luis Rodriguez Lopez-Canizares. A methodological note on the convergence of sequences of random variables. International Journal of Statistics and Probability, 1(2):35–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/18071.

Gajewski:2012:ECA

[17] Byron J. Gajewski, Stephen D. Simon, and Susan E. Carlson. On the existence of constant accrual rates in clinical trials and direction for future research. International Journal of Statistics and Probability, 1 (2):43–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/18072.

Isogai:2012:SOA

[18] Eiichi Isogai and Chikara Uno. On second-order approximations to the risk in estimating the exponential mean by a two-stage procedure. International Journal of Statistics and Probability, 1(2):47–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/18284.

Khan:2012:SIM

[19] Shahedul A. Khan, Masud Rana, Longhai Li, and Joel A. Dubin. A statistical investigation to monitor and understand atmospheric CFC decline with the spatial-longitudinal bent-cable model. International Journal of Statistics and Probability, 1(2):56–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/17253.

Rois:2012:MLB

[20] Rumana Rois, Tapati Basak, and Ajit Kumar Majumder. Marginal likelihood-based one-sided LR test for testing higher order autocorrelation in presence of nuisance parameters — a distance based approach. International Journal of Statistics and Probability, 1(2):69–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/18084.

Abushal:2012:EBG

[21] Tahani Abdulrahmaan Abushal and Areej M. Al-Zaydi. Estimation based on generalized order statistics from a mixture of two Rayleigh distributions. International Journal of Statistics and Probability, 1(2): 79–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/16088.

Adekeye:2012:MSR

[22] Kayode Samuel Adekeye. Modified simple robust control chart based on median absolute deviation. International Journal of Statistics and Probability, 1(2):91–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/17440.

Brown:2012:DCM

[23] Marlo Kay Brown. Detection of changes in a multinomial process. International Journal of Statistics and Probability, 1(2):96–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/19668.

Provost:2012:DQE

[24] Serge B. Provost and Ali Akbar Mohsenipour. The distribution of quadratic expressions in elliptically contoured vectors. International Journal of Statistics and Probability, 1(2):103–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/19300.

Ayyildiz:2012:SNR

[25] Ezgi Ayyildiz, Vilda Purutcuoğlu Gazi, and Ernst Wit. A short note on resolving singularity problems in covariance matrices. International Journal of Statistics and Probability, 1(2):113–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/18993.

Zhang:2012:RCM

[26] Jianfeng Zhang, David J. Olive, and Ping Ye. Robust covariance matrix estimation with canonical correlation analysis. International Journal of Statistics and Probability, 1(2):119–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/19093.

Cattle:2012:DQI

[27] Brian Andrew Cattle, Paul D. Baxter, Thomas J. Flemming, Christopher Peter Gale, David C. Mitchell, Mark S. Gilthorpe, Julian A. Scott, Carolyn Czoski-Murray, and Christopher McCabe. Data quality improvement, data linkage and multiple imputation in the UK National Vascular Database. International Journal of Statistics and Probability, 1 (2):137–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/17083.

Schweizer:2012:WVT

[28] Karl Schweizer. A weighted version of the tau-equivalent model of measurement for items with ordered response categories. International Journal of Statistics and Probability, 1(2):151–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/21437.

Sewell:2012:EMH

[29] Martin Victor Sewell. The efficient market hypothesis: Empirical evidence. International Journal of Statistics and Probability, 1(2):164– ??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/17911.

Suzuki:2012:HQP

[30] Joe Suzuki. The Hannan–Quinn proposition for linear regression. International Journal of Statistics and Probability, 1(2):179–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/20323.

Liang:2012:AMP

[31] Yuanyuan Liang and Keumhee C. Carriere. Allocating multiple patients at a time in multiple-objective response-adaptive repeated measurement designs. International Journal of Statistics and Probability, 1 (2):186–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/21438.

Xie:2012:SSL

[32] Yanyan Xie and Zhengping Shi. Stochastic stabilization of linear systems driven by reflecting Brownian motion. International Journal of Statistics and Probability, 1(2):198–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/20499.

Diakarya:2012:SSF

[33] Barro Diakarya, Koté Blami, and Soumaïla Moussa. Spatial stochastic framework for sampling time parametric max-stable processes. International Journal of Statistics and Probability, 1(2):203–??, ???? 2012. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/20231.

Haidara:2012:SEG

[34] Mohamed Cheikh Haidara and Gane Samb Lo. Statistical estimation of gap of decomposability of the general poverty index. International Journal of Statistics and Probability, 1(2):211–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/21752.

Erhardt:2012:BSI

[35] Erik Barry Erhardt, Balgobin Nandram, and Jai Won Choi. Bayesian simultaneous intervals for small areas: An application to variation in maps. International Journal of Statistics and Probability, 1(2):229–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/20714.

Diakarya:2012:GAC

[36] Barro Diakarya, Diouf Saliou, and Koté Blami. Geostatistical analysis with conditional extremal copulas. International Journal of Statistics and Probability, 1(2):244–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/20613.

Lo:2012:FGH

[37] Gane Samb Lo and El Hadji Deme. A functional generalized hill process and its uniform theory. International Journal of Statistics and Probability, 1(2):250–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/21754.

Ospina:2012:BCC

[38] Raydonal Ospina and Silvia L. P. Ferrari. On bias correction in a class of inflated beta regression models. International Journal of Statistics and Probability, 1(2):269–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/20694.

Prodromou:2012:SCM

[39] Theodosia Prodromou. Students’ construction of meanings about the coordination of the two epistemological perspectives on distribution. International Journal of Statistics and Probability, 1(2):283–??, ???? 2012. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/16104.

Borroni:2013:CKS

[40] Claudio Giovanni Borroni and Paola Maddalena Chiodini. A competitor of the Kolmogorov–Smirnov test for the goodness-of-fit problem. International Journal of Statistics and Probability, 2(1):1–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/20663.

Zhang:2013:SPM

[41] Shaowei Zhang. Studies on the probabilistic model for ship-bridge collisions. International Journal of Statistics and Probability, 2(1): 16–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/23258.

BenHariz:2013:EVR

[42] Samir Ben Hariz and Ali Kabui. Empirical value at risk for weak dependent random variables. International Journal of Statistics and Probability, 2(1):24–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/21077.

Lee:2013:AMM

[43] Seung-Hwan Lee, Philip Deng, and Eun-Joo Lee. Analysis of multiple myeloma life expectancy using copula. International Journal of Statistics and Probability, 2(1):44–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/20465.

Kamagate:2013:QML

[44] Amadou Kamagate and Ouagnina Hili. The quasi maximum likelihood approach to statistical inference on a nonstationary multivariate ARFIMA process. International Journal of Statistics and Probability, 2 (1):53–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/20249.

Gui:2013:MOP

[45] Wenhao Gui. A Marshall–Olkin power log-normal distribution and its applications to survival data. International Journal of Statistics and Probability, 2(1):63–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/22765.

Zenga:2013:MSL

[46] Mariangela Zenga, Marshall H. Adele, and Giordano Sabrina. Modelling students’ length of stay at university using Coxian phase-type distributions. International Journal of Statistics and Probability, 2(1): 73–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/23191.

Olive:2013:AOR

[47] David Olive. Asymptotically optimal regression prediction intervals and prediction regions for multivariate data. International Journal of Statistics and Probability, 2(1):90–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/21593.

Su:2013:GLD

[48] Steve Su, Abeer Hasan, and Wei Ning. The RS generalized lambda distribution based calibration model. International Journal of Statistics and Probability, 2(1):101–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/24144.

Yamamoto:2013:IME

[49] Kouji Yamamoto, Ryota Shinjo, and Sadao Tomizawa. Improved measure on extended marginal homogeneity for ordinal square contingency tables.

International Journal of Statistics and Probability, 2(1):108–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/21142.

Prodromou:2013:EPS

[50] Theodosia Prodromou. Estimating parameters from samples: Shuttling between spheres. International Journal of Statistics and Probability, 2 (1):113–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/23597.

Woolston:2013:MIC

[51] Andrew Woolston, Yu-Kang Tu, Mark S. Gilthorpe, and Paul D. Baxter. Measuring the impact of collinearity in epidemiological research. International Journal of Statistics and Probability, 2(2):1–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/22821.

Alola:2013:APT

[52] Andrew Adewale Alola, Mustafa Tunay, and Violet Alola. Analysis of possibility theory for reasoning under uncertainty. International Journal of Statistics and Probability, 2(2):12–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/22906.

Manoj:2013:MGB

[53] Chandrabose Manoj, Pushpa Wijekoon, and Roshan Darshana Yapa. The McDonald generalized beta-binomial distribution: A new binomial mixture distribution and simulation based comparison with its nested distributions in handling overdispersion. International Journal of Statistics and Probability, 2(2):24–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/23491.

Al-Zahrani:2013:NCP

[54] Bander Al-Zahrani, Javid Gani Dar, and Abdulkader Daghistani. A note on closure properties of classes of discrete lifetime distributions. International Journal of Statistics and Probability, 2(2):42–??, ???? 2013. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/23481.

Hernandez-Bastida:2013:BPA

[55] Agustin Hernandez-Bastida and M. Pilar Fernandez-Sanchez. The Bayes premium in an aggregate loss Poisson–Lindley model with structure function STSP. International Journal of Statistics and Probability, 2(2): 47–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/24501.

Ferland:2013:SWS

[56] Rene Ferland and François Watier. Switch-when-safe multiperiod meanvariance strategies. International Journal of Statistics and Probability, 2(2):59–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/24061.

Schweizer:2013:TFA

[57] Karl Schweizer. A threshold-free approach to the study of the structure of binary data. International Journal of Statistics and Probability, 2(2):67–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/26608.

Sanyal:2013:EUS

[58] Suman Sanyal and David Grow. Existence and uniqueness for stochastic dynamic equations. International Journal of Statistics and Probability, 2(2):77–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/25520.

Grover:2013:BAE

[59] Gurprit Grover, Alka Sabharwal, and Juhi Mittal. A Bayesian approach for estimating onset time of nephropathy for type 2 diabetic patients under various health conditions. International Journal of Statistics and Probability, 2(2):89–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/26221.

Ivanescu:2013:NBS

[60] Andrada Ivanescu. A note on bivariate smoothing for two-dimensional functional data. International Journal of Statistics and Probability, 2 (2):102–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/26157.

Nixon:2013:MHT

[61] John Harold Nixon. On multiple hypothesis testing maximizing the average power. International Journal of Statistics and Probability, 2(2): 112–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/25360.

Prodromou:2013:IIR

[62] Theodosia Prodromou. Informal inferential reasoning: Interval estimates of parameters. International Journal of Statistics and Probability, 2(2): 136–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/25979.

Sajeevkumar:2013:EPB

[63] N. K. Sajeevkumar and M. R. Irshad. Estimation of the parameters of bivariate normal distribution with equal coefficient of variation using concomitants of order statistics. International Journal of Statistics and Probability, 2(3):1–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/27795.

Alkenani:2013:PSI

[64] Ali Alkenani and Keming Yu. Penalized single-index quantile regression. International Journal of Statistics and Probability, 2(3):12–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/25490.

Alzaghal:2013:EFD

[65] Ahmad Alzaghal, Felix Famoye, and Carl Lee. Exponentiated T − X family of distributions with some applications. International Journal of Statistics and Probability, 2(3):31–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/25914.

Makri:2013:CRO

[66] Frosso S. Makri, Zaharias M. Psillakis, and Anastasios N. Arapis. Counting runs of ones with overlapping parts in binary strings ordered linearly and circularly. International Journal of Statistics and Probability, 2(3):50–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/28864.

Senger:2013:MCS

[67] Ötüken Senger. A Monte Carlo study to assess the impact of kurtosis on statistical power of Wald–Wolfowitz test. International Journal of Statistics and Probability, 2(3):61–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/24439.

Ndri:2013:HDE

[68] Aubin N’dri and Ouagnina Hili. Hellinger distance estimation of strongly dependent multi-dimensional Gaussian processes. International Journal of Statistics and Probability, 2(3):70–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/28899.

Harrison:2013:MLC

[69] Wendy J. Harrison, Mark S. Gilthorpe, Amy Downing, and Paul D. Baxter. Multilevel latent class modelling of colorectal cancer survival status at three years and socioeconomic background whilst incorporating stage of disease. International Journal of Statistics and Probability, 2 (3):85–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/28900.

El-Monsef:2013:DAS

[70] M. M. E. Abd El-Monsef and M. M. Seyam. Discriminating among several semiparametric models. International Journal of Statistics and Probability, 2(3):96–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/29000.

Satman:2013:NAD

[71] Mehmet Hakan Satman. A new algorithm for detecting outliers in linear regression. International Journal of Statistics and Probability, 2 (3):101–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/28207.

Gneyou:2013:CLT

[72] Kossi Essona Gneyou. A central limit theorem for a nonparametric maximum conditional hazard rate estimator in presence of right censoring. International Journal of Statistics and Probability, 2(3):110–??, ???? 2013.

CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/27185.

Molenberghs:2013:CBB

[73] Geert Molenberghs, Michael G. Kenward, Geert Verbeke, Samuel Iddi, and Achmad Efendi. On the connections between bridge distributions, marginalized multilevel models, and generalized linear mixed models. International Journal of Statistics and Probability, 2(4):1–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/28461.

Al-Saadony:2013:CDV

[74] Muhannad Faiz Al-Saadony, Paul Hewson, and Julian Stander. Cholesky decomposition for the Vasicek interest rate model. International Journal of Statistics and Probability, 2(4):22–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/29163.

Sinay:2013:BEF

[75] Marick S. Sinay, Chi-Wen Hsu, and John S. J. Hsu. Bayesian estimation with flexible prior for the covariance structure of linear mixed effects models. International Journal of Statistics and Probability, 2 (4):29–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/30287.

Furno:2013:QTS

[76] Marilena Furno and Domenico Vistocco. Qu test for structural breaks in quantile regressions. International Journal of Statistics and Probability, 2(4):42–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/30289.

Bialek:2013:MAR

[77] Jacek Bialek. Measuring average rate of return of pensions: A discrete, stochastic and continuous price index approaches. International Journal of Statistics and Probability, 2(4):56–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/31196.

Li:2013:AGF

[78] Yin Li and Keumhee Chough Carriere. Assessing goodness of fit of exponential random graph models. International Journal of Statistics and Probability, 2(4):64–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/27900.

Riggs:2013:PIS

[79] Kent Riggs, Dean Young, and Jeremy J. Becnel. Probability inequalities for the sum of random variables when sampling without replacement. International Journal of Statistics and Probability, 2(4):75–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/29556.

Quang:2013:RPG

[80] Phung Duy Quang. Ruin probability in a generalized risk process under interest force with homogeneous Markov chain premiums. International Journal of Statistics and Probability, 2(4):85–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/30223.

Odondi:2013:PPS

[81] Lang’o Odondi. Performance of principal stratification method adjusting for treatment noncompliance in two arms of a randomized trial. International Journal of Statistics and Probability, 2(4):93–??, ???? 2013. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/30351.

Arcede:2014:CSM

[82] Jayrold P. Arcede, Miraluna L. Herrera, Mark Jolet A. Cubelo, and Alfredo M. Grino. Chi-square mixture of transformed/inverse transformed gamma family. International Journal of Statistics and Probability, 3 (1):1–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/31148.

Elmi:2014:DMC

[83] Mohamed Elmi. Detection of multiple change points by the filtered derivative and false discovery rate. International Journal of Statistics and Probability, 3(1):12–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/32943.

Djeddour:2014:KGD

[84] K. Djeddour and S. Meddahi. Kernel gradient of density of probability estimate from contaminated associated observations. International Journal of Statistics and Probability, 3(1):24–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/32944.

Shahtahmassebi:2014:BMI

[85] Golnaz Shahtahmassebi and Rana Moyeed. Bayesian modelling of integer data using the generalised Poisson difference distribution. International Journal of Statistics and Probability, 3(1):35–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/30335.

Khalaf:2014:QSF

[86] Rehab F. Khalaf. Queueing systems with four different main server’s interruptions and a stand-by server. International Journal of Statistics and Probability, 3(1):49–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/33262.

Kuruwita:2014:TEN

[87] Chinthaka Kuruwita, Colin Gallagher, and K. B. Kulasekera. Testing equality of nonparametric quantile regression functions. International Journal of Statistics and Probability, 3(1):55–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/31998.

Satman:2014:SLR

[88] Mehmet Hakan Satman and Mustafa Altunbey. Selecting location of retail stores using artificial neural networks and Google Places API. International Journal of Statistics and Probability, 3(1):67–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/32949.

Al-Saiari:2014:MOE

[89] Arwa Y. Al-Saiari, Lamya A. Baharith, and Salwa A. Mousa. Marshall– Olkin extended Burr Type XII distribution. International Journal of Statistics and Probability, 3(1):78–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/32346.

Latif:2014:EST

[90] Sumaia A. Latif and Pedro A. Morettin. Estimation of a Spearmantype multivariate measure of local dependence. International Journal of Statistics and Probability, 3(2):1–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/30940.

Kourbatov:2014:DMP

[91] Alexei Kourbatov. The distribution of maximal prime gaps in Cramér’s probabilistic model of primes. International Journal of Statistics and Probability, 3(2):18–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/35285.

Pal:2014:MRB

[92] Baidyanath Pal and Babulal Seal. Multivariate relationships between physiologic and anthropometric variables: A data based analysis. International Journal of Statistics and Probability, 3(2):30–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/31340.

Rambaud:2014:RVF

[93] Salvador Cruz Rambaud and Antonio Luis Rodriguez Lopez-Canizares. Random variables fundamental in probability and sigma-complete convergence. International Journal of Statistics and Probability, 3(2): 46–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/35417.

Adabor:2014:ITR

[94] Emmanuel S. Adabor, George K. Acquaah-Mensah, and Francis T. Oduro. Inferring transcriptional regulatory relationships among genes in breast cancer: An application of Bayes’ Theorem. International Journal of Statistics and Probability, 3(2):52–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/33357.

Cordeiro:2014:NFD

[95] Gauss M. Cordeiro, Luis H. de Santana, Edwin M. M. Ortega, and Rodrigo R. Pescim. A new family of distributions: Libby–Novick beta. International Journal of Statistics and Probability, 3(2):63–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/28413.

Mode:2014:ESM

[96] Charles J. Mode. Estimating statistical measures of pleiotropic and epistatic effects in the genomic era. International Journal of Statistics and Probability, 3(2):81–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/35702.

Doku-Amponsah:2014:LDB

[97] K. Doku-Amponsah, F. O. Mettle, and T. Ansah-Narh. Large deviations, basic information theorem for fitness preferential attachment random networks. International Journal of Statistics and Probability, 3 (2):101–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/34289.

Doku-Amponsah:2014:EAM

[98] K. Doku-Amponsah. Exponential approximation, method of types for empirical neighbourhood distributions of random graphs by random allocations. International Journal of Statistics and Probability, 3(2): 110–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/33755.

Medeiros:2014:TIA

[99] Francisco M. C. Medeiros, Antônio H. M. da Silva-Júnior, Dione M. Valença, and Silvia L. P. Ferrari. Testing inference in accelerated failure time models. International Journal of Statistics and Probability, 3 (2):121–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/35111.

Smith:2014:RAIa

[100] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 3, no. 2. International Journal of Statistics and Probability, 3(2):132–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/36363.

Wu:2014:AEE

[101] Yu Wu, Yong Lin, Chin-Shang Li, Shou-En Lu, and Weichung Joe Shih. Asymptotic efficiency of an exponential cure model when cure information is partially known. International Journal of Statistics and Probability, 3(3):1–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/34438.

Arrabal:2014:BAU

[102] Claude Thiago Arrabal, Marinho Gomes de Andrade Filho, and Karina Paula dos Santos Silva. Bayesian approach using latent variable for zero truncated Poisson distribution: Application for species-area relationship. International Journal of Statistics and Probability, 3(3): 18–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/35902.

Elmi:2014:POF

[103] Mohamed Elmi. The parameters optimization of filtered derivative for change points analysis. International Journal of Statistics and Probability, 3(3):29–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/37770.

Yuan:2014:ANM

[104] Ao Yuan, Yizheng Wei, Kepher Makambi, and Hongkun Wang. Approximate nonparametric maximum likelihood estimation for interval censoring model Case II. International Journal of Statistics and Probability, 3(3):44–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/36130.

Marie:2014:PFO

[105] Nicolas Marie. A pathwise fractional one compartment intra-veinous bolus model. International Journal of Statistics and Probability, 3(3): 65–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/38179.

Al-Dayian:2014:SIS

[106] G. R. Al-Dayian, A. A. El-Helbawy, and H. R. Rezk. Statistical inference for a simple constant stress model based on censored sampling data from the Kumaraswamy Weibull distribution. International Journal of Statistics and Probability, 3(3):80–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/38180.

Montoya:2014:SSC

[107] Eduardo L. Montoya, Nehemias Ulloa, and Victoria Miller. A simulation study comparing knot selection methods with equally spaced knots in a penalized regression spline. International Journal of Statistics and Probability, 3(3):96–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/36992.

Wu:2014:MMM

[108] Longyang Wu and Richard J. Cook. Marginal methods for multivariate failure times under event-dependent censoring. International Journal of Statistics and Probability, 3(3):111–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/37529.

Chen:2014:MEC

[109] Feng Chen and Kani Chen. Modeling event clustering using the mmemory Cox-type self-exciting intensity model. International Journal of Statistics and Probability, 3(3):126–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/38984.

Gharib:2014:SNC

[110] M. Gharib, M. M. Ramadan, and Kh. A. H. Al-Ajmi. Some new characterizations of Markov–Bernoulli geometric distribution related to random sums. International Journal of Statistics and Probability, 3(3): 138–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/38009.

Moghimipour:2014:CDT

[111] Ida Moghimipour and Malihe Ebrahimpour. Comparing decision tree method over three data mining software. International Journal of Statistics and Probability, 3(3):147–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/37872.

Smith:2014:RAIb

[112] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 3, no. 3. International Journal of Statistics and Probability, 3(3):157–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/38987.

Yurachkivsky:2014:CMS

[113] Andriy Yurachkivsky. A conditional mean square estimate for the solution of a SDE. International Journal of Statistics and Probability, 3 (4):1–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/39453.

Acquah:2014:TSB

[114] Stella B. Acquah, Richard K. Avuglah, and Emmanuel Harris. Time series behavior of the volume of wood products export in Ghana. International Journal of Statistics and Probability, 3(4):12–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/37986.

Mahnken:2014:EVU

[115] Jonathan D. Mahnken, Xueyi Chen, Alexandra R. Brown, Eric D. Vidoni, Sandra A. Billinger, and Byron J. Gajewski. Evaluating variables as unbiased proxies for other measures: Assessing the step test exercise prescription as a proxy for the maximal, high-intensity peak oxygen consumption in older adults. International Journal of Statistics and Probability, 3(4):25–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/38088.

Niang:2014:TBN

[116] Mouhamed Amine Niang, Guy Martial Nkiet, and Aliou Diop. Testing bivariate normality based on nonlinear canonical analysis. International Journal of Statistics and Probability, 3(4):35–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/40592.

Saeed:2014:MEG

[117] Bashiru I. I. Saeed, S. K. Appiah, N. N. N. Nsowah Nuamah, Louis Munyakazi, and A. A. I. Musah. Model equivalence in general linear models: Set-to-zero, sum-to-zero restrictions, and extra sum of squares method. International Journal of Statistics and Probability, 3(4):42–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/41186.

Iwundu:2014:MQC

[118] M. P. Iwundu and D. W. Ebong. Modified quick convergent inflow algorithm for solving linear programming problems. International Journal of Statistics and Probability, 3(4):54–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/39933.

Iwundu:2014:BOE

[119] M. P. Iwundu and E. B. Albert-Udochukwuka. On the behaviour of Doptimal exact designs under changing regression polynomials. International Journal of Statistics and Probability, 3(4):67–??, ???? 2014. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/39990.

Smith:2014:RAIc

[120] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 3, no. 4. International Journal of Statistics and Probability, 3(4):86–??, ???? 2014. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/41821.

Provost:2015:PAS

[121] Serge B. Provost and Susan Z. Sheng. Polynomially adjusted saddlepoint density approximations. International Journal of Statistics and Probability, 4(1):1–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/41014.

He:2015:EAL

[122] Nie He, Yuan-Biao Zhang, Zhen Zhang, and Xin-Guang Lv. Evaluation of academic level of sci-tech journals based on rough set and TOPSIS. International Journal of Statistics and Probability, 4(1):12–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/40684.

Dai:2015:ESF

[123] Hongsheng Dai. Exact simulation for fork-join networks with heterogeneous service. International Journal of Statistics and Probability, 4 (1):19–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/41295.

Tao:2015:DFM

[124] Aiyang Tao, Yong Lin, Jose Pinheiro, and Weichung Joe Shih. Dose finding method in joint modeling of efficacy and safety endpoints in Phase II studies. International Journal of Statistics and Probability, 4 (1):33–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/41443.

Marra:2015:FBB

[125] Giampiero Marra and Rosalba Radice. Flexible bivariate binary models for estimating the efficacy of phototherapy for newborns with jaundice. International Journal of Statistics and Probability, 4(1):46–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/41892.

Gharib:2015:CSN

[126] M. Gharib, M. M. Mohie El-Din, and A. Sharawy. Characterization of the skew-normal distribution via order statistics and record values. International Journal of Statistics and Probability, 4(1):59–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/42102.

Hu:2015:ESP

[127] Yunjiao Hu, Guangqiang Lan, and Chong Zhang. The explicit solution and precise distribution of CKLS model under Girsanov transform. International Journal of Statistics and Probability, 4(1):68–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/42640.

Doku-Amponsah:2015:LDP

[128] K. Doku-Amponsah, F. O. Mettle, and E. N. N. Nortey. Large deviation principle for the empirical degree measure of preferential attachment random graphs. International Journal of Statistics and Probability, 4(1):76–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/35771.

Young:2015:IDP

[129] Phil Dean Young. Independence distribution-preserving covariance structures for the likelihood ratio test for LXB = 0 in the general linear model. International Journal of Statistics and Probability, 4(1):87–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/41049.

Doku-Amponsah:2015:LDR

[130] Kwabena Doku-Amponsah. Large deviation result for the empirical locality measure of typed random geometric graphs. International Journal of Statistics and Probability, 4(1):94–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/41819.

Meng:2015:CBC

[131] Si-Jing Meng, Yuan-Biao Zhang, Gui-Sen Xu, and Jun-Ta Wu. College basketball coach evaluation based on dynamic DHGF evaluation model. International Journal of Statistics and Probability, 4(1):100–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/44196.

Ramos:2015:BXN

[132] Manoel Wallace A. Ramos, Ana Percontini, Gauss M. Cordeiro, and Ronaldo V. da Silva. The Burr XII negative binomial distribution with applications to lifetime data. International Journal of Statistics and Probability, 4(1):109–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/44352.

Olive:2015:VTM

[133] David J. Olive, Lasanthi C. R. Pelawa Watagoda, and Hasthika S. Rupasinghe Arachchige Don. Visualizing and testing the multivariate linear regression model. International Journal of Statistics and Probability, 4 (1):126–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/43460.

Salha:2015:RNW

[134] Raid B. Salha and Hazem I. El Shekh Ahmed. Reweighted Nadaraya– Watson estimator of the regression mean. International Journal of Statistics and Probability, 4(1):138–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/39379.

Sharma:2015:EEV

[135] Dinesh Sharma, Amanda Miller, and Caroline Hollingsworth. Estimating explained variation of a latent scale dependent variable underlying a binary indicator of event occurrence. International Journal of Statistics and Probability, 4(1):148–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/39718.

Sinha:2015:RAI

[136] Sanjoy K. Sinha and Abdus Sattar. A robust approach to identifying differential circulating miRNAs in breast cancer. International Journal of Statistics and Probability, 4(1):155–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/44670.

Yi:2015:RFD

[137] Changsheng Yi and Qiumei Jin. Random fuzzy decision models for pharmaceutical R&D project investment under uncertainty. International Journal of Statistics and Probability, 4(1):170–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/43984.

Smith:2015:RAIa

[138] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 4, no. 1. International Journal of Statistics and Probability, 4(1):182–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/44767.

Changpetch:2015:MSP

[139] Pannapa Changpetch and Dennis K. J. Lin. Model selection for Poisson regression via association rules analysis. International Journal of Statistics and Probability, 4(2):1–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/46231.

Xia:2015:BMM

[140] Michelle Xia and David P. M. Scollnik. A Bayesian mixture model accounting for zeros and negatives in the loss triangle. International Journal of Statistics and Probability, 4(2):10–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/46233.

Al-Zahrani:2015:SPR

[141] Bander Al-Zahrani and Mashail Al-Sobhi. On some properties of the reversed variance residual lifetime. International Journal of Statistics and Probability, 4(2):24–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/43232.

Sarhan:2015:ETP

[142] Ammar M. Sarhan, Bruce Smith, and David C. Hamilton. Estimation of P (Y < X) for a two-parameter bathtub shaped failure rate distribution. International Journal of Statistics and Probability, 4(2):33–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/46723.

Gregoire:2015:NRI

[143] Timothy G. Gregoire. Nonparallel regressions with indicator variables. International Journal of Statistics and Probability, 4(2):46–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/45155.

Umelo-Ibemere:2015:QPP

[144] Ngozi C. Umelo-Ibemere and Harrison O. Amuji. Quantile plots of the prediction variance for partially replicated central composite design. International Journal of Statistics and Probability, 4(2):55–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/43972.

Oluyede:2015:IAW

[145] Broderick O. Oluyede, Macaulay Okwuokenye, and Karl E. Peace. Inequalities and approximations of weighted distributions by Lindley reliability measures, and the Lindley–Cox model with applications. International Journal of Statistics and Probability, 4(2):61–??, ???? 2015. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/44836.

Aboueissa:2015:CTM

[146] Abou El-Makarim A. Aboueissa. Comparison of two means of two lognormal distributions when data is singly censored. International Journal of Statistics and Probability, 4(2):73–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/47291.

Ashoori:2015:MGE

[147] Farnoosh Ashoori, Malihe Ebrahimpour, Atena Gharib, and Abolghasem Bozorgnia. Minimal generalized extreme value distribution and its application in modeling of minimum temperature. International Journal of Statistics and Probability, 4(2):87–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/47805.

Iwundu:2015:CSQ

[148] M. P. Iwundu and I. E. Ndiyo. Comparative study of the Quick Convergent Inflow Algorithm (QCIA) and the Modified Quick Convergent Inflow Algorithm (MQCIA). International Journal of Statistics and Probability, 4(2):95–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/46244.

Ahsanullah:2015:CWX

[149] M. Ahsanullah, A. Alzaatreh, I. Ghosh, and G. G. Hamedani. Characterizations of the Weibull–X and Burr XII negative binomial families of distributions. International Journal of Statistics and Probability, 4 (2):113–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/46474.

Koh:2015:ARI

[150] Daniel Koh. Assessing relative importance using RSP scoring to generate Variable Importance Factor (VIF). International Journal of Statistics and Probability, 4(2):123–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/46446.

Hu:2015:BPG

[151] Kejia Hu, Jaesik Choi, Alex Sim, and Jiming Jiang. Best predictive generalized linear mixed model with predictive lasso for high-speed network data analysis. International Journal of Statistics and Probability, 4(2):132–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/48070.

Smith:2015:RAIb

[152] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 4, no. 2. International Journal of Statistics and Probability, 4(2):149–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/48230.

Khan:2015:MML

[153] Shahedul Khan. Modeling a mixture of linear and changepoint trajectories for longitudinal time-series data. International Journal of Statistics and Probability, 4(3):1–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/46121.

Almheidat:2015:SGF

[154] Maalee Almheidat, Felix Famoye, and Carl Lee. Some generalized families of Weibull distribution: Properties and applications. International Journal of Statistics and Probability, 4(3):18–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/47078.

Yu:2015:GLR

[155] Han Yu. Generalized likelihood ratio tests based on the asymptotic variant of the minimax approach. International Journal of Statistics and Probability, 4(3):36–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/47092.

Albassam:2015:CPD

[156] Mohammed Albassam. Characterizations of probability distributions with completely monotone hazard functions. International Journal of Statistics and Probability, 4(3):55–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/47364.

Aldahmani:2015:UEL

[157] Saeed Aldahmani and Hongsheng Dai. Unbiased estimation for linear regression when n < v. International Journal of Statistics and Probability, 4(3):61–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/47855.

Kuruwita:2015:RSI

[158] Chinthaka Kuruwita. Regularized single index quantile regression model. International Journal of Statistics and Probability, 4(3):74–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/48894.

Nanthakumar:2015:DHU

[159] A. Nanthakumar, S. Kanbur, and E. Wilson. On the detection of heteroscedasticity by using CUSUM range distribution. International Journal of Statistics and Probability, 4(3):88–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/50691.

Hamedani:2015:KHC

[160] G. G. Hamedani and I. Ghosh. Kumaraswamy-half-Cauchy distribution: Characterizations and related results. International Journal of Statistics and Probability, 4(3):94–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/50692.

Phillips:2015:TSE

[161] Jeremy Phillips, Steven B. Caudill, and Franklin G. Mixon. Tournament seeding efficiency and home court advantage: College Basketball’s National Invitation Tournament. International Journal of Statistics and Probability, 4(3):101–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/47409.

Chowdhury:2015:SMM

[162] Salehin K. Chowdhury and Sanjoy K. Sinha. Semiparametric marginal models for binary longitudinal data. International Journal of Statistics and Probability, 4(3):107–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/48546.

Ahsanullah:2015:CDC

[163] M. Ahsanullah, G. G. Hamedani, and M. Maadooliat. Characterizations of distributions via conditional expectation of generalized order statistics. International Journal of Statistics and Probability, 4(3):121–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/49073.

Aboueissa:2015:CMT

[164] Abou El-Makarim A. Aboueissa. Comparison of means of two log-normal distributions when data is multiply censored. International Journal of Statistics and Probability, 4(3):127–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/50977.

Chaudhuri:2015:PMC

[165] G Chaudhuri. On periodic maintenance of a coherent system. International Journal of Statistics and Probability, 4(3):145–??, ???? 2015. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/51057.

Bibi:2015:NSC

[166] Abdelouahab Bibi and Fateh Merahi. A note on L_{2}-structure of continuous-time bilinear processes with time-varying coefficients. International Journal of Statistics and Probability, 4(3):150–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/51242.

Weissman-Miller:2015:NSD

[167] Deborah Weissman-Miller and Kay C. Graham. Novel scale development for fear of falling and falls: Analysed using a semiparametric ratio estimator (SPRE). International Journal of Statistics and Probability, 4 (3):161–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/50127.

Yi:2015:MSD

[168] Changsheng Yi and Qiumei Jin. A multi-stages decision approach for managerial flexibility of energy R&D project under fuzzy environment. International Journal of Statistics and Probability, 4(3):169–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/51350.

Prodromou:2015:SER

[169] Theodosia Prodromou. Students’ emerging reasoning about data tables of large-scale data. International Journal of Statistics and Probability, 4(3):181–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/50246.

Smith:2015:RAIc

[170] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 4, no. 3. International Journal of Statistics and Probability, 4(3):198–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/51658.

Iwundu:2015:OPR

[171] M. P. Iwundu. Optimal partially replicated cube, star and center runs in face-centered central composite designs. International Journal of Statistics and Probability, 4(4):1–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/51587.

daSilva:2015:SMA

[172] Alan Ricardo da Silva and Iracema Veiga Madeira Mauriz. A SAS macro for adaptive spatial sampling. International Journal of Statistics and Probability, 4(4):20–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/51693.

Schweizer:2015:DCF

[173] Karl Schweizer, Xuezhu Ren, Tengfei Wang, and Florian Zeller. Does the constraint of factor loadings impair model fit and accuracy in parameter estimation? International Journal of Statistics and Probability, 4(4):40–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/53390.

Douglas:2015:INP

[174] Richard Douglas. Imaginary number probability in Bayesian-type inference. International Journal of Statistics and Probability, 4(4):51– ??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/53391.

Hossain:2015:ETR

[175] Shakhawat Hossain and Hatem A. Howlader. Estimation techniques for regression model with zero-inflated Poisson data. International Journal of Statistics and Probability, 4(4):64–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/53393.

Sarhan:2015:ITP

[176] Ammar M. Sarhan and Joseph Apaloo. Inferences for a two-parameter lifetime distribution with bathtub shaped hazard based on censored data. International Journal of Statistics and Probability, 4(4):77–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/53898.

Asiedu:2015:SEF

[177] Louis Asiedu, Atinuke O. Adebanji, Francis Oduro, and Felix O. Mettle. Statistical evaluation of face recognition techniques under variable environmental constraints. International Journal of Statistics and Probability, 4(4):93–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/53899.

daSilva:2015:EBX

[178] Ronaldo V. da Silva, Frank Gomes-Silva, Manoel Wallace A. Ramos, and Gauss M. Cordeiro. The exponentiated Burr XII Poisson distribution with application to lifetime data. International Journal of Statistics and Probability, 4(4):112–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/52145.

Afify:2015:TMO

[179] Ahmed Z. Afify, G. G. Hamedani, Indranil Ghosh, and M. E. Mead. The transmuted Marshall–Olkin Fréchet distribution: Properties and applications. International Journal of Statistics and Probability, 4(4): 132–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/54370.

Smith:2015:RAId

[180] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 4, no. 4. International Journal of Statistics and Probability, 4(4):149–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/54372.

Cordeiro:2016:MXL

[181] Gauss M. Cordeiro, Edwin M. M. Ortega, G. G. Hamedani, and Diogo A. Garcia. The McBurr XII and log-McBurr XII models with applications to lifetime data. International Journal of Statistics and Probability, 5 (1):1–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/51846.

Brimacombe:2016:NLS

[182] Michael Brimacombe. A note on linear and second order significance testing in nonlinear models. International Journal of Statistics and Probability, 5(1):19–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/53152.

Ajao:2016:TDM

[183] Isaac O. Ajao, Femi J. Ayoola, and Joseph O. Iyaniwura. Temporal disaggregation methods in flow variables of economic data: Comparison study. International Journal of Statistics and Probability, 5(1):36–??, ???? 2016.

CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/55020.

Wei:2016:PCR

[184] Chuanhua Wei and Xiaonan Wang. Principal components regression estimation in semiparametric partially linear additive models. International Journal of Statistics and Probability, 5(1):46–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/54629.

Roberts:2016:SCP

[185] Gareth O. Roberts and Jeffrey S. Rosenthal. Surprising convergence properties of some simple Gibbs samplers under various scans. International Journal of Statistics and Probability, 5(1):51–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/55570.

Kadhem:2016:RDI

[186] Safaa K. Kadhem, Paul Hewson, and Irene Kaimi. Recursive deviance information criterion for the hidden Markov model. International Journal of Statistics and Probability, 5(1):61–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/53814.

Salha:2016:AKE

[187] Raid B. Salha, Hazem I. El Shekh Ahmed, and Hossam O. El-Sayed. Adaptive kernel estimation of the conditional quantiles. International Journal of Statistics and Probability, 5(1):79–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/54764.

Naka:2016:ADC

[188] Mayumi Naka and Ritei Shibata. Asymptotic distribution of Cramér– von Mises statistic when contamination exists. International Journal of Statistics and Probability, 5(1):90–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/54776.

Iwundu:2016:RII

[189] M. P. Iwundu and S. U. Ekpo. Rules for identifying the initial design points for use in the quick convergent inflow algorithm. International Journal of Statistics and Probability, 5(1):98–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/55148.

Brobbey:2016:WZI

[190] Anita Brobbey, Aerambamoorthy Thavaneswaran, and Saumen Mandal. Wrapped zero-inflated Poisson distribution and its properties. International Journal of Statistics and Probability, 5(1):111–??, ???? 2016. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/55891.

Smith:2016:RAIa

[191] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 5, no. 1. International Journal of Statistics and Probability, 5(1):116–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/56134.

Luo:2016:CAD

[192] Yao Luo and Eunji Lim. On consistency of absolute deviations estimators of convex functions. International Journal of Statistics and Probability, 5(2):1–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/56476.

Weissman-Miller:2016:PSP

[193] Deborah Weissman-Miller. On predicting survival in prostate cancer: Using an extended maximum spacing method at the change point of the semiparametric ratio estimator (SPRE). International Journal of Statistics and Probability, 5(2):19–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/57236.

Arfi:2016:CRK

[194] Mounir Arfi. On the convergence rate for a kernel estimate of the regression function. International Journal of Statistics and Probability, 5 (2):29–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/57237.

Hovda:2016:TDE

[195] Sigve Hovda. Transmetric density estimation. International Journal of Statistics and Probability, 5(2):35–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/57238.

White:2016:EAN

[196] Alexander K. White and Samir K. Safi. The efficiency of artificial neural networks for forecasting in the presence of autocorrelated disturbances. International Journal of Statistics and Probability, 5(2):51–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/56670.

Liang:2016:RDR

[197] You Liang, A. Thavaneswaran, and B. Abraham. Recent developments in recursive estimation for time series models. International Journal of Statistics and Probability, 5(2):59–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/57314.

Nofal:2016:KTE

[198] Zohdy M. Nofal, Ahmed Z. Afify, Haitham M. Yousof, Daniele C. T. Granzotto, and Francisco Louzada. Kumaraswamy transmuted exponentiated additive Weibull distribution. International Journal of Statistics and Probability, 5(2):78–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/56452.

Xu:2016:MME

[199] Yingmei Xu, Kane Ladji, and Diawara Daouda. Mixed method of extreme value theory, with application to the calculation of the portion of each claim payable by the reinsurer of excess of sinister. International Journal of Statistics and Probability, 5(2):100–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/57638.

Smith:2016:RAIb

[200] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, Vol. 5, No. 2. International Journal of Statistics and Probability, 5(2):109–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/57854.

Thavaneswaran:2016:EWZ

[201] Aerambamoorthy Thavaneswaran, Saumen Mandal, and Dharini Pathmanathan. Estimation for wrapped zero inflated Poisson and wrapped Poisson distributions. International Journal of Statistics and Probability, 5(3):1–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/56307.

Hashimoto:2016:NFR

[202] Elizabeth M. Hashimoto, Gauss M. Cordeiro, Edwin M. M. Ortega, and G. G. Hamedani. New flexible regression models generated by gamma random variables with censored data. International Journal of Statistics and Probability, 5(3):9–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/56529.

Bar-Lev:2016:KRE

[203] Shaul K. Bar-Lev, Benzion Boukai, and Zinoviy Landsman. The Kendal– Ressel exponential dispersion model: Some statistical aspects and estimation. International Journal of Statistics and Probability, 5(3): 32–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/58260.

Erdogan:2016:LTE

[204] Birsen Eygi Erdogan. Long-term examination of bank crashes using panel logistic regression: Turkish banks failure case. International Journal of Statistics and Probability, 5(3):42–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/58837.

Tahir:2016:BSE

[205] Mohamed Tahir. Bayesian sequential estimation of the inverse of the Pareto shape parameter. International Journal of Statistics and Probability, 5(3):49–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/59001.

Kanwai:2016:UAV

[206] M. E. Kanwai, O. E. Asiribo, and A. Isah. Use of auxiliary variables and asymptotically optimum estimators in double sampling. International Journal of Statistics and Probability, 5(3):55–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/56508.

Hovda:2016:PTD

[207] Sigve Hovda. Properties of transmetric density estimation. International Journal of Statistics and Probability, 5(3):63–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/55493.

Baker:2016:MBD

[208] Rose D. Baker and Ian G. McHale. Making better decisions: Can minimizing frequentist risk help? International Journal of Statistics and Probability, 5(3):80–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/59005.

Chalise:2016:TSE

[209] Prabhakar Chalise, Eric Chicken, and Daniel McGee. Time scales in epidemiological analysis: An empirical comparison. International Journal of Statistics and Probability, 5(3):91–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/58234.

Guan:2016:SLL

[210] Li Guan and Ying Wan. A strong law of large numbers for set-valued negatively dependent random variables. International Journal of Statistics and Probability, 5(3):102–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/58143.

Smith:2016:RAIc

[211] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, Vol. 5, No. 3. International Journal of Statistics and Probability, 5(3):113–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/59433.

Al-Zahrani:2016:ERW

[212] Bander Al-Zahrani. On the estimation of reliability of weighted Weibull distribution: A comparative study. International Journal of Statistics and Probability, 5(4):1–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/59454.

Carneiro:2016:GLR

[213] Hérica P. A. Carneiro and Dione M. Valença. Gradient and likelihood ratio tests in cure rate models. International Journal of Statistics and Probability, 5(4):9–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/58419.

Iwundu:2016:ASO

[214] Mary Paschal Iwundu. Alternative second-order n-point spherical response surface methodology design and their efficiencies. International Journal of Statistics and Probability, 5(4):22–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/58860.

Yamamoto:2016:CDS

[215] Kouji Yamamoto, Yuya Matsuda, and Sadao Tomizawa. Collapsed double symmetry model and its decomposition for square contingency tables. International Journal of Statistics and Probability, 5(4):31–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/57386.

Iki:2016:DEC

[216] Kiyotaka Iki, Akira Shibuya, and Sadao Tomizawa. Diagonal exponent conditional symmetry model for square contingency tables with ordered categories. International Journal of Statistics and Probability, 5 (4):38–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/57598.

Korzeniowski:2016:DRC

[217] Andrzej Korzeniowski and Rachel Traylor. Dynamic reliability of a cluster server. International Journal of Statistics and Probability, 5 (4):45–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/60558.

Gollob:2016:PPC

[218] Michael Gollob, Jeffrey S. Rosenthal, and Kevin Thorpe. The probability of pathogenicity in clinical genetic testing: A solution for the variant of uncertain significance. International Journal of Statistics and Probability, 5(4):52–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/60559.

Angelini:2016:DBP

[219] Pierpaolo Angelini and Antonio Maturo. On the dichotomy between the psychological or empirical aspect of subjective probability and the logical or geometrical one. International Journal of Statistics and Probability, 5(4):59–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/60560.

Nanthakumar:2016:MKT

[220] Ampalavanar Nanthakumar. Multivariate kurtosis as a tool for comparing copula models. International Journal of Statistics and Probability, 5(4):67–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/58463.

Tschirk:2016:PID

[221] Wolfgang Tschirk. The principle of indifference does not lead to contradictions. International Journal of Statistics and Probability, 5(4): 79–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/59617.

Hyodo:2016:ECB

[222] Yoshifumi Hyodo, Masahide Kuwada, and Hiromu Yumiba. On the existence conditions for balanced fractional 2^{m} factorial designs of resolution R^{∗}({1}|Ω_{ℓ}) with N < ν_{ℓ}(m). International Journal of Statistics and Probability, 5(4):84–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/56892.

Mukhoti:2016:MCH

[223] Sujay Mukhoti and Pritam Ranjan. Mean-correction and higher order moments for a stochastic volatility model with correlated errors. International Journal of Statistics and Probability, 5(4):102–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/60912.

Hesse:2016:EAR

[224] Christian A. Hesse, John B. Ofosu, and Samuel K. Darkwah. The effect of age on road traffic fatality index in Ghana. International Journal of Statistics and Probability, 5(4):111–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/60834.

Guillaume:2016:CSP

[225] Tristan Guillaume. Computation of the survival probability of Brownian motion with drift when the absorbing boundary is a piecewise affine or piecewise exponential function of time. International Journal of Statistics and Probability, 5(4):119–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/60989.

Alzaghal:2016:SWP

[226] Ahmad Alzaghal, Indranil Ghosh, and Ayman Alzaatreh. On shifted Weibull–Pareto distribution. International Journal of Statistics and Probability, 5(4):139–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/60990.

Lim:2016:ESF

[227] Eunji Lim and Mina Attallah. Estimation of smooth functions via convex programs. International Journal of Statistics and Probability, 5(4): 150–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/59900.

Smith:2016:RAId

[228] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 5, no. 4. International Journal of Statistics and Probability, 5(4):156–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/61024.

Maiti:2016:TPF

[229] Saran Ishika Maiti. Tau-proportionate form of congeneric covariance structure. International Journal of Statistics and Probability, 5(5): 1–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/59844.

Elgmati:2016:QME

[230] Entisar A. Elgmati and Nadia B. Gregni. Quartile method estimation of two-parameter exponential distribution data with outliers. International Journal of Statistics and Probability, 5(5):12–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/62129.

Zhao:2016:CBS

[231] Guolong Zhao. Considering both statistical and clinical significance. International Journal of Statistics and Probability, 5(5):16–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/62130.

Rahman:2016:AEP

[232] Azizur Rahman, Junbin Gao, Catherine D’Este, and Syed Ejaz Ahmed. An assessment of the effects of prior distributions on the Bayesian predictive inference. International Journal of Statistics and Probability, 5 (5):31–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/59513.

Wu:2016:ECP

[233] Yanhong Wu. Estimation of change-point and post-change parameters after adaptive sequential CUSUM test in an exponential family. International Journal of Statistics and Probability, 5(5):43–??, ???? 2016. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/61399.

Schweizer:2016:CFC

[234] Karl Schweizer, Xuezhu Ren, and Michael Schreiner. Construct-focused configural invariance for measures showing a multi-dimensional structure and application to exchange test data. International Journal of Statistics and Probability, 5(5):55–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/62630.

Smith:2016:RAIe

[235] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 5, no. 5. International Journal of Statistics and Probability, 5(5):67–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/62648.

Fenga:2016:TSC

[236] Livio Fenga. Time series chaos detection and assessment via scale dependent Lyapunov exponent. International Journal of Statistics and Probability, 5(6):1–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/61955.

Depren:2016:DSV

[237] Serpil Kilic Depren and Özer Depren. Determination of support vector boundaries in generalized maximum entropy for multilevel models. International Journal of Statistics and Probability, 5(6):10–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/62123.

Gobbi:2016:CBU

[238] Fabio Gobbi. Convolution based unit root processes: a simulation approach. International Journal of Statistics and Probability, 5(6): 22–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/63734.

Appiah:2016:LHD

[239] E. A. Appiah and G. S. Ladde. Linear hybrid deterministic dynamic modeling for time-to-event processes: State and parameter estimations. International Journal of Statistics and Probability, 5(6):32–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/63735.

Mai:2016:IBP

[240] Kristi Mai and Qingyang Zhang. Identification of biomarkers for predicting the overall survival of ovarian cancer patients: a sparse group lasso approach. International Journal of Statistics and Probability, 5 (6):57–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/63932.

Iwok:2016:SMF

[241] Iberedem A. Iwok. Seasonal modelling of Fourier series with linear trend. International Journal of Statistics and Probability, 5(6):65–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/61797.

Urge:2016:ETW

[242] Birhanu Worku Urge, Kepher Makambi, and Anthony Wanjoya. Estimating three-way latent interaction effects in structural equation modeling. International Journal of Statistics and Probability, 5(6):73–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/62832.

Joshua:2016:EDS

[243] Ogunyiola Ayorinde Joshua, Peter N. Mwita, and Carolyn N. Njenga. Estimating dependence structure and risk of financial market crash. International Journal of Statistics and Probability, 5(6):85–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/62980.

Beauducel:2016:RET

[244] André Beauducel, Christopher Harms, and Norbert Hilger. Reliability estimates for three factor score estimators. International Journal of Statistics and Probability, 5(6):94–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/63934.

Al-Hussain:2016:CAD

[245] Raed Al-Hussain and Reza Khorramshahgol. A comprehensive analysis of the determinants of swap problem in the supply chain of the petroleum industry. International Journal of Statistics and Probability, 5(6):108–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/63985.

Hesse:2016:ARR

[246] Christian A. Hesse, Francis T. Oduro, John B. Ofosu, and Emmanuel D. Kpeglo. Assessing the risk of road traffic fatalities across sub-populations of a given geographical zone, using a modified Smeed’s model. International Journal of Statistics and Probability, 5(6):121–??, ???? 2016. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/64109.

Smith:2016:RAIf

[247] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 5, no. 6. International Journal of Statistics and Probability, 5(6):133–??, ???? 2016. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/64110.

Cortese:2017:EAU

[248] Giuliana Cortese. Estimating the area under the ROC curve with modified profile likelihoods. International Journal of Statistics and Probability, 6(1):1–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/64476.

Ji:2017:SLP

[249] Chunlin Ji. On sequential learning for parameter estimation in particle algorithms for state-space models. International Journal of Statistics and Probability, 6(1):13–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/56968.

Yurachkivsky:2017:LDG

[250] Andriy Yurachkivsky. Limit distribution of a generalized Ornstein– Uhlenbeck process. International Journal of Statistics and Probability, 6(1):24–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/65098.

Guan:2017:LTN

[251] Li Guan. Limit theorems for negatively dependent fuzzy set-valued random variables. International Journal of Statistics and Probability, 6 (1):40–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/65099.

Askin:2017:PES

[252] Oykum Esra Askin, Deniz Inan, and Ali Hakan Buyuklu. Parameter estimation of shared frailty models based on particle swarm optimization. International Journal of Statistics and Probability, 6(1):48–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/65100.

Allab:2017:FPT

[253] Imene Allab and François Watier. First-passage time estimation of diffusion processes through time-varying boundaries with an application in finance. International Journal of Statistics and Probability, 6 (1):59–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/63088.

Ladji:2017:EVT

[254] Kane Ladji, Diawara Daouda, and Diallo Moumouni. Extreme value theory: a new characterization of the distribution function for the mixed method. International Journal of Statistics and Probability, 6 (1):71–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/65235.

VonTress:2017:ISM

[255] Mark Von Tress. Inverse sampling for McNemar’s test. International Journal of Statistics and Probability, 6(1):78–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/65236.

Kolev:2017:CEV

[256] Nikolai Kolev and Jayme Pinto. Characterizations of extreme value extended Marshall–Olkin models with exponential marginals. International Journal of Statistics and Probability, 6(1):87–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/65237.

Jiahui:2017:IDT

[257] Tang Jiahui, Zhang Yuanbiao, Peng Churu, and Huang Xinxin. Installation and dispatch of the traffic patrol service platform. International Journal of Statistics and Probability, 6(1):95–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/65238.

Brimacombe:2017:DBS

[258] Michael B. Brimacombe. A directional Bayesian significance test for equality of variances. International Journal of Statistics and Probability, 6 (1):102–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/65239.

Granzotto:2017:SSM

[259] Daniele Cristina Tita Granzotto, Josmar Mazucheli, and Francisco Louzada. Statistical study of monthly rainfall trends by using the transmuted power Lindley distribution. International Journal of Statistics and Probability, 6(1):111–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/57080.

Aryal:2017:TLG

[260] Gokarna R. Aryal, Edwin M. Ortega, G. G. Hamedani, and Haitham M. Yousof. The Topp–Leone generated Weibull distribution: Regression model, characterizations and applications. International Journal of Statistics and Probability, 6(1):126–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/65369.

Zhang:2017:SMC

[261] Qiang Zhang and Michael R. Kosorok. Statistical monitoring of clinical trials using Brownian bridges. International Journal of Statistics and Probability, 6(1):142–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/65370.

Smith:2017:RAIa

[262] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 6, no. 1. International Journal of Statistics and Probability, 6(1):152–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/65397.

Iwok:2017:JWT

[263] Iberedem A. Iwok. Justification of Wold’s theorem and the unbiasedness of a stable vector autoregressive time series model forecasts. International Journal of Statistics and Probability, 6(2):1–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/64392.

Wen:2017:EIT

[264] Jialin Wen, Min Zou, Yikai Ma, and Hao Luo. Evaluating the influence of taxi subsidy programs on mitigating difficulty getting a taxi in basis of taxi empty-loaded rate. International Journal of Statistics and Probability, 6(2):9–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/66355.

Bouza:2017:RTE

[265] Carlos N. Bouza, Amer Ibrahim Al-Omari, Agustín Santiago, and Jose M. Sautto. Ratio type estimation using the knowledge of the auxiliary variable for ranking and estimating. International Journal of Statistics and Probability, 6(2):21–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/47050.

Lee:2017:MCD

[266] Eun-Joo Lee, Noah Klumpe, Jonathan Vlk, and Seung-Hwan Lee. Modeling conditional dependence of stock returns using a copula-based GARCH model. International Journal of Statistics and Probability, 6 (2):32–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/66356.

Andersson:2017:CAI

[267] Per Gösta Andersson. A classroom approach to illustrate transformation and bootstrap confidence interval techniques using the Poisson distribution. International Journal of Statistics and Probability, 6(2): 42–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/63242.

Lamberti:2017:ITD

[268] Giuseppe Lamberti and Tomas Aluja Banet. Invariance test: Detecting difference between latent variables structure in partial least squares path modeling. International Journal of Statistics and Probability, 6 (2):54–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/66429.

Tarsitano:2017:MAD

[269] Agostino Tarsitano and Ilaria L. Amerise. A model to approximate the distribution of rank order associations. International Journal of Statistics and Probability, 6(2):66–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/66509.

Dey:2017:DDP

[270] Sanku Dey, Bander Al-Zahrani, and Samerah Basloom. Dagum distribution: Properties and different methods of estimation. International Journal of Statistics and Probability, 6(2):74–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/64605.

Traylor:2017:RCT

[271] Rachel Lunde Traylor and Andrzej Korzeniowski. Reliability of a clustered-task server under modulated correlation. International Journal of Statistics and Probability, 6(2):93–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/65562.

Yu:2017:TIC

[272] Yuan Yu, Dhiman Bhadra, and Balgobin Nandram. Tests of independence for a 2 × 2 contingency table with random margins. International Journal of Statistics and Probability, 6(2):106–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/66542.

Granzotto:2017:OBS

[273] Daniele Cristina Tita Granzotto, Vera Lucia Damasceno Tomazella, and Francisco Louzada. Objective Bayesian snalysis for the complementary exponential geometric model applied to cancer data. International Journal of Statistics and Probability, 6(2):122–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/65025.

Lawal:2017:SMM

[274] Bayo H. Lawal. On some mixture models for over-dispersed binary data. International Journal of Statistics and Probability, 6(2):134–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/66623.

Smith:2017:RAIb

[275] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 6, no. 2. International Journal of Statistics and Probability, 6(2):148–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/66627.

Lim:2017:EMS

[276] Eunji Lim and Fanny Gonzalez. Estimation of multivariate smooth functions via convex programs. International Journal of Statistics and Probability, 6(3):1–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/65724.

NjamenNjomen:2017:CNA

[277] Didier Alain Njamen Njomen. Convergence of the Nelson–Aalen estimator in competing risks. International Journal of Statistics and Probability, 6(3):9–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/67019.

Cordeiro:2017:EGS

[278] Gauss M. Cordeiro, Thiago A. N. de Andrade, Marcelo Bourguignon, and Frank Sinatra Gomes da Silva. The exponentiated generalized standardized half-logistic distribution. International Journal of Statistics and Probability, 6(3):24–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/67120.

Kolev:2017:DME

[279] Nikolai Kolev and Jayme Pinto. Dependence modeling in energy markets using Sibuya-type copulas. International Journal of Statistics and Probability, 6(3):43–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/67291.

Zeng:2017:GVP

[280] Bilin Zeng, Kang Chen, and Cong Wang. Geometric views of partial correlation coefficient in regression analysis. International Journal of Statistics and Probability, 6(3):51–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/66936.

Carrasco:2017:EKD

[281] Jalmar M. F. Carrasco and Gauss M. Cordeiro. An extension of the Kumaraswamy distribution. International Journal of Statistics and Probability, 6(3):61–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/64575.

Santana:2017:GEW

[282] Tiago V. F. Santana, Edwin M. M. Ortega, Gauss M. Cordeiro, and Adriano K. Suzuki. Generalized exponentiated Weibull linear model in the presence of covariates. International Journal of Statistics and Probability, 6(3):75–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/62506.

Wen:2017:EMR

[283] Jialin Wen and Min Zou. Evaluating model of road traffic in open housing estate based on cellular automaton. International Journal of Statistics and Probability, 6(3):99–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/67574.

Saghir:2017:WDB

[284] Aamir Saghir, G. G. Hamedani, Sadaf Tazeem, and Aneeqa Khadim. Weighted distributions: A brief review, perspective and characterizations. International Journal of Statistics and Probability, 6(3):109–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/67733.

Gbohounme:2017:SSR

[285] Idelphonse Leandre Tawanou Gbohounme, Oscar Owino Ngesa, and Jude Eggoh. Self-selecting robust logistic regression model. International Journal of Statistics and Probability, 6(3):132–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/67734.

deAndrade:2017:GGP

[286] Thiago A. N. de Andrade, Luz Milena Zea Fernandez, Frank Gomes-Silva, and Gauss M. Cordeiro. The gamma generalized Pareto distribution with applications in survival analysis. International Journal of Statistics and Probability, 6(3):141–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/67750.

Barnabani:2017:PTD

[287] Marco Barnabani. A parametric test to discriminate between a linear regression model and a linear latent growth model. International Journal of Statistics and Probability, 6(3):157–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/67751.

Scalise:2017:HMM

[288] Kathleen Scalise. Hybrid measurement models for technology-enhanced assessments through mIRT–Bayes. International Journal of Statistics and Probability, 6(3):168–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/67182.

Kamari:2017:CBC

[289] Osman Kamari. Characterizations based on cumulative entropy of the last-order statistics. International Journal of Statistics and Probability, 6(3):183–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/67849.

Emmanuel:2017:EDR

[290] Oti-Boateng Emmanuel. Estimating disease risk of diabetes cases in the presence of underreporting. International Journal of Statistics and Probability, 6(3):188–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/67850.

Wilcox:2017:LCB

[291] Rand R. Wilcox. Linear contrasts based on an extension of the Wilcoxon– Mann–Whitney approach. International Journal of Statistics and Probability, 6(3):198–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/67851.

Almendra-Arao:2017:NMF

[292] Félix Almendra-Arao, Hortensia J. Reyes-Cervantes, and José J. Castro-Alva. A new margin function for anti-infective trials. International Journal of Statistics and Probability, 6(3):204–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/67967.

Smith:2017:RAIc

[293] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 6, no. 3. International Journal of Statistics and Probability, 6(3):213–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/67976.

Okorie:2017:ALL

[294] I. E. Okorie, A. C. Akpanta, J. Ohakwe, D. C. Chikezie, C. U. Onyemachi, and E. O. Obi. The adjusted log-logistic generalized exponential distribution with application to lifetime data. International Journal of Statistics and Probability, 6(4):1–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/68740.

Ma:2017:CFM

[295] Jishan Ma and Yuanbiao Zhang. Commodity future money flows trading strategies based on HMM. International Journal of Statistics and Probability, 6(4):16–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/68741.

Rikimaru:2017:NIS

[296] Yuuki Rikimaru and Ritei Shibata. Non-identifiability of simultaneous spatial autoregressive model and singularity of Fisher information matrix. International Journal of Statistics and Probability, 6(4):31–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/68085.

Yusuf:2017:UQR

[297] Oyindamola B. Yusuf, Ayo S. Adebowale, Elijah A. Bamgboye, Temitayo Odusote, Olusanmi Iyabode, and Oladapo A. Ladipo. Use of quantile regression in determining factors associated with BMI among vulnerable adolescents in Rivers State, Nigeria. International Journal of Statistics and Probability, 6(4):39–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/69050.

Bortolini:2017:EGG

[298] Juliano Bortolini, Marcelino A. R. Pascoa, Renato Ribeiro de Lima, and Anderson C. S. Oliveira. The extended generalized gamma geometric distribution. International Journal of Statistics and Probability, 6(4):48–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/69051.

Farsipour:2017:ECM

[299] N. Sanjari Farsipour and A. Asgharzadeh. Estimating the common mean of k normal populations with known variance. International Journal of Statistics and Probability, 6(4):70–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/69088.

Fadonougbo:2017:JAS

[300] Renaud Fadonougbo and George O. Orwa. Jump adapted scheme of a non mark dependent jump diffusion process with application to the Merton jump diffusion model. International Journal of Statistics and Probability, 6(4):80–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/69087.

Smith:2017:RAId

[301] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, Vol. 6, No. 4. International Journal of Statistics and Probability, 6(4):98–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/69131.

Najarzadegan:2017:TWM

[302] Hossein Najarzadegan, Mohammad Hossein Alamatsaz, and Saied Hayati. Truncated Weibull-G more flexible and more reliable than beta-G distribution. International Journal of Statistics and Probability, 6 (5):1–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/68488.

Lim:2017:NTC

[303] Eunji Lim and Elizabeth Tavarez. Nonparametric tests for convexity/ monotonicity/positivity of multivariate functions with noisy observations. International Journal of Statistics and Probability, 6(5):18–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/68727.

Hsu:2017:IMR

[304] John S. J. Hsu and Thomas Leonard. Insurance mortality rates, performance indicators, and possibly monotonic population proportions. International Journal of Statistics and Probability, 6(5):29–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/69596.

Temesgen:2017:APM

[305] Aboma Temesgen. Application of Poisson mixed combined models for identifying correlations of CD4 count progression in HIV infected TB patients during ART treatment period. International Journal of Statistics and Probability, 6(5):42–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/67856.

Zareamoghaddam:2017:LDE

[306] Hossein Zareamoghaddam, Serge B. Provost, and Ejaz Ahmed. A logdensity estimation methodology applicable to massive bivariate data. International Journal of Statistics and Probability, 6(5):53–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/69827.

Hassan:2017:GAW

[307] Amal S. Hassan, Saeed E. Hemeda, Sudhansu S. Maiti, and Sukanta Pramanik. The generalized additive Weibull-G family of distributions. International Journal of Statistics and Probability, 6(5):65–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/69828.

Otunuga:2017:TVP

[308] Olusegun Michael Otunuga. Time varying parameter estimation scheme for a linear stochastic differential equation. International Journal of Statistics and Probability, 6(5):84–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/69440.

Groenitz:2017:IEA

[309] Heiko Groenitz. Improving estimation accuracy in nonrandomized response questioning methods by multiple answers. International Journal of Statistics and Probability, 6(5):101–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/70003.

Zurbenko:2017:AKZ

[310] Igor G. Zurbenko and Mingzeng Sun. Applying Kolmogorov–Zurbenko adaptive R-software. International Journal of Statistics and Probability, 6(5):110–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/69632.

Lawal:2017:CNS

[311] Bayo H. Lawal. Correcting for non-sum to 1 estimated probabilities in applications of discrete probability models to count data. International Journal of Statistics and Probability, 6(5):119–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/70180.

Ahsanullah:2017:SCE

[312] M. Ahsanullah and M. Z. Anis. Some characterizations of exponential distribution. International Journal of Statistics and Probability, 6 (5):132–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/70181.

Prodromou:2017:MBI

[313] Theodosia Prodromou. Model-based informal inference. International Journal of Statistics and Probability, 6(5):140–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/70247.

Smith:2017:RAIe

[314] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, Vol. 6, No. 5. International Journal of Statistics and Probability, 6(5):148–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/70278.

Sha:2017:BIS

[315] Naijun Sha and Hao Yang Teng. A Bayes inference for step-stress accelerated life testing. International Journal of Statistics and Probability, 6(6):1–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/69742.

Asai:2017:ATE

[316] Manabu Asai and Michael McAleer. Asymptotic theory for extended asymmetric multivariate GARCH processes. International Journal of Statistics and Probability, 6(6):13–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/69933.

Jensen:2017:UHO

[317] D. R. Jensen and D. E. Ramirez. Use of Hotelling’s T^{2}: Outlier diagnostics in mixtures. International Journal of Statistics and Probability, 6(6):24–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/70608.

Schweizer:2017:CVL

[318] Karl Schweizer, Stefan Troche, and Siegbert Reiß. Can variances of latent variables be scaled in such a way that they correspond to eigenvalues? International Journal of Statistics and Probability, 6(6):35–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/70609.

Morin:2017:CTQ

[319] Samantha Morin, Myron Hlynka, and Shan Xu. Comments on a two queue network. International Journal of Statistics and Probability, 6 (6):43–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/69772.

Lewis:2017:RAF

[320] Frank D. Lewis and Gordon J. Horn. Rasch analysis and functional measurement in post-hospital brain injury rehabilitation. International Journal of Statistics and Probability, 6(6):50–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/70735.

Mokhlis:2017:IES

[321] Nahed A. Mokhlis, Emad J. Ibrahim, and Dina M. Gharieb. Interval estimation of stress-strength reliability for a general exponential form distribution with different unknown parameters. International Journal of Statistics and Probability, 6(6):60–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/70737.

Gharib:2017:MOE

[322] M. Gharib, B. I. Mohammed, and W. E. R. Aghel. Marshll–Olkin extended inverse Pareto distribution and its application. International Journal of Statistics and Probability, 6(6):71–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/70738.

Hasanthika:2017:ACA

[323] Hasanthika, N. H. E., Jayasekara, and L. A. L. W. Analyzing the customer attrition using survival techniques. International Journal of Statistics and Probability, 6(6):85–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/69932.

Bae:2017:HTC

[324] Taehan Bae and Jingjiao Chen. On heavy-tailed crack distribution for loss severity modeling. International Journal of Statistics and Probability, 6(6):92–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/70501.

Akpan:2017:DMA

[325] Emmanuel Alphonsus Akpan and Imoh Udo Moffat. Detection and modeling of asymmetric GARCH effects in a discrete-time series. International Journal of Statistics and Probability, 6(6):111–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/70857.

Shu:2017:NML

[326] Di Shu and Wenqing He. A new method for logistic model assessment. International Journal of Statistics and Probability, 6(6):120–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/71134.

Herranz:2017:URT

[327] Ed Herranz, James Gentle, and George Wang. Unit roots in time series with changepoints. International Journal of Statistics and Probability, 6 (6):127–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/71167.

Assonken:2017:FNL

[328] Patrick Assonken and Gangaram S. Ladde. A family of non linear models in a market with semi Markov regimes: Application to the commodity and the derivative market. International Journal of Statistics and Probability, 6(6):137–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/71298.

Mutter:2017:USA

[329] Louis Mutter and Steven B. Kim. Using simple alternative hypothesis to increase statistical power in sparse categorical data. International Journal of Statistics and Probability, 6(6):158–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/71300.

Eissa:2017:EKW

[330] Fathy Helmy Eissa. The exponentiated Kumaraswamy–Weibull distribution with application to real data. International Journal of Statistics and Probability, 6(6):167–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/70606.

Smith:2017:RAIf

[331] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, Vol. 6, No. 6. International Journal of Statistics and Probability, 6(6):183–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/71477.

Rahmouni:2018:EGT

[332] Mohieddine Rahmouni and Ayman Orabi. The Exponential-Generalized Truncated Geometric (EGTG) distribution: A new lifetime distribution. International Journal of Statistics and Probability, 7(1):1–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/70950.

DallAglio:2018:CBP

[333] Marco Dall’Aglio and Theodore P. Hill. Characterization of the Bayesian posterior distribution in terms of self-information. International Journal of Statistics and Probability, 7(1):21–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/71011.

Tse:2018:ALT

[334] Siu Keung Tse and Chang Ding. Accelerated life test sampling plans under progressive Type II interval censoring with random removals. International Journal of Statistics and Probability, 7(1):26–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/70820.

Hamedani:2018:CIDa

[335] G. G. Hamedani. Characterizations and infinite divisibility of certain recently introduced distributions III. International Journal of Statistics and Probability, 7(1):39–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/71753.

Baharith:2018:PSE

[336] Lamya A. Baharith. Parametric and semiparametric estimations of bivariate truncated Type I generalized logistic models driven from copulas. International Journal of Statistics and Probability, 7(1):72–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/71852.

Ghitany:2018:SCR

[337] M. E. Ghitany, Ramesh C. Gupta, and Shaochen Wang. Some characterization results by conditional expectations and their applications to Lindley-type distributions. International Journal of Statistics and Probability, 7(1):86–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/71042.

Demissie:2018:DHT

[338] Million Wesenu Demissie. Determinants and heterogeneity of time-to-recovery from obstetric fistula patients; comparison of acceleration failure time and parametric shared frailty models. International Journal of Statistics and Probability, 7(1):99–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/71530.

Kitaoka:2018:SDR

[339] Yoshiyuki Kitaoka. Statistical distribution of roots of a polynomial modulo primes III. International Journal of Statistics and Probability, 7 (1):115–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/72465.

Patterson:2018:MMM

[340] Joshua Patterson and Andrzej Korzeniowski. M/M/1 model with unreliable service. International Journal of Statistics and Probability, 7 (1):125–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/72701.

Smith:2018:RAIa

[341] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, Vol. 7, No. 1. International Journal of Statistics and Probability, 7(1):137–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/72703.

Granzotto:2018:TWR

[342] Daniele Cristina Tita Granzotto, Carlos Aparecido dos Santos, and Francisco Louzada. The transmuted Weibull regression model: an application to Type 2 Diabetes Mellitus data. International Journal of Statistics and Probability, 7(2):1–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/71886.

Makubate:2018:BWF

[343] Boikanyo Makubate, Broderick O. Oluyede, Gofaone Motobetso, Shujiao Huang, and Adeniyi F. Fagbamigbe. The beta Weibull-G family of distributions: Model, properties and application. International Journal of Statistics and Probability, 7(2):12–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/73031.

Boubakar:2018:SCI

[344] Traoré Boubakar, Diabaté Lassina, Touré Belco, and Fané Abdou. The shortest confidence interval for the mean of a normal distribution. International Journal of Statistics and Probability, 7(2):33–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/73032.

Liu:2018:WNP

[345] Xiaoyang Liu, Hengyang Liu, Ya Luo, and Chao Liu. WSN node positioning and mathematics modeling based on genetic method. International Journal of Statistics and Probability, 7(2):39–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/73033.

Angelini:2018:GEN

[346] Pierpaolo Angelini and Angela De Sanctis. On a geometric extension of the notion of exchangeability referring to random events. International Journal of Statistics and Probability, 7(2):50–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/73330.

Hossain:2018:PDB

[347] M. J. Hossain and A. K. Majumder. Proposed distance-based test for testing multivariate multiple regression coefficients under restricted alternatives. International Journal of Statistics and Probability, 7(2): 56–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/72779.

Tashkandy:2018:BBG

[348] Yusra A. Tashkandy, Maha A. Omair, and Abdulhamid A. Alzaid. Bivariate and bilateral gamma distributions. International Journal of Statistics and Probability, 7(2):66–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/73589.

Lorio:2018:DES

[349] Jennifer Lorio, Norou Diawara, and Lance A. Waller. Density estimation of spatio-temporal point patterns using Moran’s statistic. International Journal of Statistics and Probability, 7(2):80–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/73728.

Congdon:2018:AIM

[350] Peter Congdon. Assessing impacts on mortality of lifestyle factors: Allowing for model uncertainty. International Journal of Statistics and Probability, 7(2):91–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/73822.

Smith:2018:RAIb

[351] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 7, no. 2. International Journal of Statistics and Probability, 7(2):108–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/73823.

Baffoun:2018:ENS

[352] Hatem Baffoun, Mekki Hajlaoui, and Abdeljelil Farhat. On estimating non-standard discrete distributions using adaptive MCMC methods. International Journal of Statistics and Probability, 7(3):1–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/73058.

B:2018:MBR

[353] B and Adekanmbi. Multiple binomial regression models of learning style preferences of students of Sidhu School, Wilkes University. International Journal of Statistics and Probability, 7(3):9–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/74096.

Yusuf:2018:MEZ

[354] Oyindamola Yusuf, Rotimi Afolabi, and Ayoola Ayoola. Modelling excess zeros in count data with application to antenatal care utilisation. International Journal of Statistics and Probability, 7(3):22–??, ???? 2018.

CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/72850.

Hamedani:2018:CIDb

[355] G. G. Hamedani. Characterizations and infinite divisibility of certain recently introduced distributions IV. International Journal of Statistics and Probability, 7(3):36–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/74860.

Zhang:2018:SRE

[356] Chenyang Zhang, Chuanhua Wei, and Bailing An. Stochastic restricted estimation in partially linear measurement error models. International Journal of Statistics and Probability, 7(3):66–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/73793.

Oluyede:2018:NGL

[357] Broderick O. Oluyede, Huybrechts F. Bindele, Boikanyo Makubate, and Shujiao Huang. A new generalized log-logistic and modified Weibull distribution with applications. International Journal of Statistics and Probability, 7(3):72–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/74861.

Yan:2018:ITC

[358] Hao Yan. The influence of the two-child policy on China’s population projection. International Journal of Statistics and Probability, 7 (3):94–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/74862.

Iki:2018:DPI

[359] Kiyotaka Iki, Shun Sato, and Sadao Tomizawa. Decomposition of parsimonious independence model using Pearson, Kendall and Spearman’s correlations for two-way contingency tables. International Journal of Statistics and Probability, 7(3):105–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/73979.

Ghosh:2018:CVB

[360] D. K. Ghosh, Shreya Ghosh, and Narenrda C. Sinohjia. On a class of variance balanced designs. International Journal of Statistics and Probability, 7(3):112–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/75049.

Smith:2018:RAIc

[361] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 7, no. 3. International Journal of Statistics and Probability, 7(3):121–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/75050.

Kim:2018:ELP

[362] Steven Kim. Explaining Lord’s Paradox in introductory statistical theory courses. International Journal of Statistics and Probability, 7(4): 1–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/75051.

Singh:2018:ATM

[363] Shatrunjai Singh, Swagata Karkare, Sudhir Baswan, and Vijendra Singh. The application of text mining algorithms in summarizing trends in antiepileptic drug research. International Journal of Statistics and Probability, 7(4):11–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/75273.

Saigusa:2018:EMH

[364] Yusuke Saigusa, Tomohisa Maruyama, Kouji Tahata, and Sadao Tomizawa. Extended marginal homogeneity model based on complementary log–log transform for square tables. International Journal of Statistics and Probability, 7(4):27–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/75639.

Karunarathna:2018:MTT

[365] K. Karunarathna, S. Brindha, and P. Paramadevan. Modeling trend in telecommunication in Sri Lanka: A case study on Internet and cellular connections. International Journal of Statistics and Probability, 7 (4):32–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/75640.

Le:2018:NBP

[366] Giang Truong Le. New bounds on Poisson approximation for random sums of independent binomial random variables. International Journal of Statistics and Probability, 7(4):43–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/74733.

Beauducel:2018:OAT

[367] André Beauducel and Norbert Hilger. On optimal allocation of treatment/ condition variance in principal component analysis. International Journal of Statistics and Probability, 7(4):50–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/76019.

Almamy:2018:TPO

[368] Jehan. A. Almamy, Mohamed Ibrahim, M. S. Eliwa, Saeed Al-mualim, and Haitham M. Yousof. The two-parameter odd Lindley Weibull lifetime model with properties and applications. International Journal of Statistics and Probability, 7(4):57–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/76020.

Doustmohammadi:2018:ISV

[369] Mehrnaz Doustmohammadi, Niloufar Shirani Bidabadi, Sumalatha Kesaveraddy, and Michael Anderson. The impact of sidewalks on vehiclepedestrian crash severity. International Journal of Statistics and Probability, 7(4):69–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/76022.

Eissa:2018:SSR

[370] Fathy Helmy Eissa. Stress-strength reliability model with the exponentiated Weibull distribution: Inferences and applications. International Journal of Statistics and Probability, 7(4):78–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/75088.

Lee:2018:BFM

[371] Tze-San Lee. The Bayes factor for the misclassified categorical data. International Journal of Statistics and Probability, 7(4):91–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/76198.

Kikechi:2018:CLP

[372] Conlet Biketi Kikechi and Richard Onyino Simwa. On comparison of local polynomial regression estimators for P = 0 and P = 1 in a model based framework. International Journal of Statistics and Probability, 7 (4):104–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/75438.

Smith:2018:RAId

[373] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 7, no. 4. International Journal of Statistics and Probability, 7(4):115–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/76199.

Bashir:2018:CRR

[374] Shakila Bashir and Mujahid Rasul. A characterization and recurrence relations of moments of the size-biased power function distribution by lower record values. International Journal of Statistics and Probability, 7(5):1–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/76152.

Zhang:2018:ASM

[375] Ru Zhang, Zi ang Lin, Shaozhen Chen, Min Zhao, and Mingjie Yuan. Adaboost–SVM multi-factor stock selection model based on Adaboost enhancement. International Journal of Statistics and Probability, 7(5): 9–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/76173.

Islam:2018:ODF

[376] Masudul Islam, Afroza Akhtar, Sirajum Munira, Md. Salauddin Khan, and Md Monzur Murshed. Optimized Dickey–Fuller test refines sign and boundary problems compare to traditional Dickey–Fuller test. International Journal of Statistics and Probability, 7(5):19–??, ???? 2018. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/76744.

Berzofsky:2018:TVG

[377] Marcus E. Berzofsky and Paul P. Biemer. Time varying grouping variables in Markov latent class analysis: Some problems and solutions. International Journal of Statistics and Probability, 7(5):28–??, ???? 2018.

CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/76745.

Kesavareddy:2018:CCS

[378] Sumalatha Kesavareddy, Kirolos Haleem, Mehrnaz Doustmohammadi, and Michael Anderson. Comparison of crash severity risk factors at signalized and stop-controlled intersections in urban and rural areas in Alabama. International Journal of Statistics and Probability, 7(5):50–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:

//www.ccsenet.org/journal/index.php/ijsp/article/view/76754.

Davis:2018:IFR

[379] Rebecca Davis, Elvis K. Donkoh, Bernard Mawah, and Blessed Amonoo. Internal financial risk management in microfinance companies: A case study of Akuapem Rural Bank, Ghana. International Journal of Statistics and Probability, 7(5):64–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/76755.

Abouelmagd:2018:LBH

[380] T. H. M. Abouelmagd. The logarithmic Burr–Hatke exponential distribution for modeling reliability and medical data. International Journal of Statistics and Probability, 7(5):73–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/76756.

Deng:2018:TSM

[381] Bo Deng and Keumhee Chough Carriere. Testing simultaneous marginal homogeneity for clustered matched-pair multinomial data. International Journal of Statistics and Probability, 7(5):86–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/76757.

P:2018:CMC

[382] M. P. and Iwundu. Construction of modified central composite designs for non-standard models. International Journal of Statistics and Probability, 7(5):95–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/76758.

Abouelmagd:2018:NFV

[383] T. H. M. Abouelmagd. A new flexible version of the Lomax distribution with applications. International Journal of Statistics and Probability, 7 (5):120–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/76826.

Smith:2018:RAIe

[384] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 7, no. 5. International Journal of Statistics and Probability, 7(5):133–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/76827.

Lawal:2018:GSC

[385] Bayo H. Lawal. GLM for some class of com-Poisson distributions with applications. International Journal of Statistics and Probability, 7(6): 1–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/76789.

Laghari:2018:SAC

[386] A. H. Laghari, M. G. H. Talpur, R. M. I. Rajput, S. J. Leghari, U. Ali, G. M. Laghari, and M. A. Sanjrani. Statistical analysis of cotton cultivated area, production and price. International Journal of Statistics and Probability, 7(6):14–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/76790.

Singh:2018:UML

[387] Shatrunjai P. Singh, Swagata Karkare, Sudhir M. Baswan, and Vijendra P. Singh. Unsupervised machine learning for co/multimorbidity analysis. International Journal of Statistics and Probability, 7(6):23– ??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/36857.

Marzjarani:2018:HMS

[388] Morteza Marzjarani. Heteroscedasticity and model selection via partitioning in fisheries data. International Journal of Statistics and Probability, 7(6):33–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/36858.

Makubate:2018:BLL

[389] Boikanyo Makubate, Broderick O. Oluyede, Neo Dingalo, and Adeniyi Francis Fagbamigbe. The beta log-logistic Weibull distribution: Model, properties and application. International Journal of Statistics and Probability, 7(6):49–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/36859.

Schweizer:2018:ASI

[390] Karl Schweizer, Siegbert Reiß, and Stefan Troche. Approaches for structural investigations of binary data using confirmatory factor models. International Journal of Statistics and Probability, 7(6):68–??, ???? 2018. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/36860.

Fang:2018:EPA

[391] Fang Fang, Yong Lin, Weichung Joe Shih, Shou-En Lu, and Guangrui Zhu. Evaluation of performance of adaptive designs based on treatment effect intervals. International Journal of Statistics and Probability, 7 (6):81–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/36892.

Nwaigwe:2018:MCM

[392] Chrysogonus C. Nwaigwe, Chukwudi J. Ogbonna, and Ojochekpa Achem. On the modeling of carbon monoxide flaring in Nigeria. International Journal of Statistics and Probability, 7(6):94–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/37168.

Soro:2018:APW

[393] Brahima Soro, Ouagnina Hili, and Youssouf Diagana. Asymptotic properties for weighted averages of longitudinal dependent data. International Journal of Statistics and Probability, 7(6):100–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/37169.

Islam:2018:EPP

[394] Khairul Islam and Tanweer Shapla. Estimation of the Poisson parameter with moment generating method. International Journal of Statistics and Probability, 7(6):113–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/37170.

Don:2018:RBO

[395] Hasthika S. Rupasinghe Arachchige Don. A relationship between the one-way MANOVA test statistic and the Hotelling Lawley trace test statistic. International Journal of Statistics and Probability, 7(6):124– ??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/37171.

Alturk:2018:CSN

[396] Lutfiah Ismail Al turk and Wejdan Saleem Al ahmadi. Comparative study of the non-homogeneous Poisson process Type-I generalized halflogistic distribution. International Journal of Statistics and Probability, 7(6):132–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/37229.

Alkan:2018:FSR

[397] Altug Alkan and Orhan Ozgur Aybar. On a family of sequences related to prime counting function. International Journal of Statistics and Probability, 7(6):149–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/37240.

Belu:2018:CDM

[398] Alexandru Belu, Wesley Maddox, and Wojbor A. Woyczynski. Copulas and dependency measures for multivariate Linnik distributions. International Journal of Statistics and Probability, 7(6):154–??, ???? 2018. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/37305.

Smith:2018:RAIf

[399] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 7, no. 6. International Journal of Statistics and Probability, 7(6):172–??, ???? 2018. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/37306.

Korir:2019:PIE

[400] Betty C. Korir and Josphat K. Kinyanjui. Parametric interval estimation of the Geeta distribution. International Journal of Statistics and Probability, 8(1):1–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/37485.

Amerise:2019:NMD

[401] Ilaria Lucrezia Amerise and Agostino Tarsitano. A new method to detect outliers in high-frequency time series. International Journal of Statistics and Probability, 8(1):16–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/37486.

Yang:2019:EMP

[402] Aotian Yang and Wanjie Sun. An exact method for power calculation for a three-arm clinical endpoint bioequivalence study. International Journal of Statistics and Probability, 8(1):25–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/37490.

Watagoda:2019:SMT

[403] Lasanthi C. R. Pelawa Watagoda. A sub-model theorem for ordinary least squares. International Journal of Statistics and Probability, 8(1): 40–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/37493.

Alturk:2019:CAN

[404] Lutfiah Ismail Al turk. Characteristics and application of the NHPP loglogistic reliability model. International Journal of Statistics and Probability, 8(1):44–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/37740.

Hamedani:2019:CCR

[405] G. G. Hamedani. Characterizations of certain recently introduced discrete distributions. International Journal of Statistics and Probability, 8(1):56–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/37823.

Mashabe:2019:NCG

[406] Baitshephi Mashabe, Broderick O. Oluyede, Adeniyi F. Fagbamigbe, Boikanyo Makubate, and Syamala Krishnannair. A new class of generalized Weibull-G family of distributions: Theory, properties and applications. International Journal of Statistics and Probability, 8(1): 73–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/37827.

Bhatti:2019:CRT

[407] Fiaz Ahmad Bhatti, G. G. Hamedani, Wenhui Sheng, and Munir Ahmad. Cubic rank transmuted modified Burr III Pareto distribution: Development, properties, characterizations and applications. International Journal of Statistics and Probability, 8(1):94–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/37829.

Nanthakumar:2019:ERC

[408] A. Nanthakumar. On the error rate comparison of the quadratic discriminant function, Euclidean distance classifier, Fisher’s linear discriminant function and the vine copulas. International Journal of Statistics and Probability, 8(1):113–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/37916.

Kim:2019:STR

[409] Steven B. Kim and Jeffrey O. Wand. Sequential testing in reliability and validity studies with repeated measurements per subject. International Journal of Statistics and Probability, 8(1):120–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/37917.

Yuan:2019:ELI

[410] Mingao Yuan and Yue Zhang. Empirical likelihood inference for partial functional linear regression models based on B-spline. International Journal of Statistics and Probability, 8(1):135–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/37918.

Doustmohammadi:2019:BRM

[411] Mehrnaz Doustmohammadi and Michael Anderson. A Bayesian regression model for estimating average daily traffic volumes for low volume roadways. International Journal of Statistics and Probability, 8 (1):143–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/37943.

Smith:2019:RAIa

[412] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 8, no. 1. International Journal of Statistics and Probability, 8(1):150–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/37947.

Patterson:2019:MMM

[413] Joshua Patterson and Andrzej Korzeniowski. M/M/1 model with unreliable service and a working vacation. International Journal of Statistics and Probability, 8(2):1–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/38162.

Moon:2019:BCD

[414] Wooyoung M. Moon and Jun S. Song. Boosted convolutional decision trees for translationally invariant pattern recognition and transfer learning. International Journal of Statistics and Probability, 8(2): 11–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/38163.

He:2019:CLM

[415] Wenqing He, Wantao Wang, Lihua Yue, and Xinhua Liu. A composite likelihood method for the analysis of multivariate survival data: An application to a PBDE study. International Journal of Statistics and Probability, 8(2):23–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/38164.

Mwanyekange:2019:BJM

[416] Josua Mwanyekange, Samuel Musili Mwalili, and Oscar Ngesa. Bayesian joint models for longitudinal and multi-state survival data. International Journal of Statistics and Probability, 8(2):34–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/38165.

Almamy:2019:EPL

[417] Jehan. A. Almamy. Extended Poisson-log-logistic distribution. International Journal of Statistics and Probability, 8(2):56–??, ???? 2019. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/38166.

Korkmaz:2019:OPL

[418] Mustafa C. Korkmaz, Emrah Altun, Haitham M. Yousof, and G. G. Hamedani. The odd power Lindley generator of probability distributions: Properties, characterizations and regression modeling. International Journal of Statistics and Probability, 8(2):70–??, ???? 2019. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/38322.

Lin:2019:MPM

[419] Tsung-Wu Lin and Hejie Lin. Mechanical proof of the Maxwell speed distribution. International Journal of Statistics and Probability, 8(2): 90–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/38323.

Mode:2019:RMU

[420] Charles J. Mode and Candace K. Sleeman. A review of the methodologies used in the derivation of formulas for parametric survival functions with illustrative numerical examples. International Journal of Statistics and Probability, 8(2):99–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/38353.

Eissa:2019:EPE

[421] Fathy H. Eissa, Shuo-Jye Wu, and Hamid H. Ahmed. Estimation of the parameters and expected test time of exponentiated Weibull lifetimes under Type II progressive censoring scheme with random removals. International Journal of Statistics and Probability, 8(2):124–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/38481.

Saigusa:2019:MDL

[422] Yusuke Saigusa, Mitsuhiro Takami, Aki Ishii, and Sadao Tomizawa. Measure of departure from local symmetry for square contingency tables. International Journal of Statistics and Probability, 8(2):140– ??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/38482.

Al-mualim:2019:EPI

[423] Saeed Al-mualim. Extended Poisson inverse Weibull distribution: Theoretical and computational aspects. International Journal of Statistics and Probability, 8(2):146–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/38507.

Marzjarani:2019:IAH

[424] Morteza Marzjarani. Iterative approaches to handling heteroscedasticity with partially known error variances. International Journal of Statistics and Probability, 8(2):159–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/38531.

Liu:2019:CNP

[425] Gang Liu, Qiang Fu, and Junjun Kang. Cavitation and negative pressure: A flexible water model molecular dynamics simulation. International Journal of Statistics and Probability, 8(2):172–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/38532.

Orumie:2019:PCA

[426] Ukamaka Cynthia Orumie and Ogbonna Onyinyechi. Principal component analysis and its derivation from singular value decomposition. International Journal of Statistics and Probability, 8(2):183–??, ???? 2019. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/38533.

Smith:2019:RAIb

[427] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 8, no. 2. International Journal of Statistics and Probability, 8(2):192–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/38637.

Kang:2019:MBD

[428] John Sang Jin Kang, Serge B. Provost, and Jiandong Ren. Momentbased density approximation techniques as applied to heavy-tailed distributions. International Journal of Statistics and Probability, 8(3): 1–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/39024.

Ghatak:2019:IBS

[429] Debolina Ghatak and Bimal K. Roy. An improved bound for security in an identity disclosure problem. International Journal of Statistics and Probability, 8(3):24–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/39033.

Marwa:2019:CTM

[430] Yohana Maiga Marwa, Isambi Sailon Mbalawata, and Samuel Mwalili. Continuous time Markov chain model for cholera epidemic transmission dynamics. International Journal of Statistics and Probability, 8 (3):32–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/39172.

Pierpaolo:2019:EAB

[431] Angelini Pierpaolo and Angela De Sanctis. An essential analogy between coherent previsions of random gains and well-behaved preferences. International Journal of Statistics and Probability, 8(3):46–??, ???? 2019. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/39173.

Bar-Lev:2019:MCM

[432] Shaul K. Bar-Lev and Ad Ridder. Monte Carlo methods for insurance risk computation. International Journal of Statistics and Probability, 8 (3):54–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/39212.

Rezgui:2019:PIN

[433] Anis Rezgui. Polynomial interpolation of normal conditional expectation. International Journal of Statistics and Probability, 8(3):75– ??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/39215.

Cheruiyot:2019:BCN

[434] Langat Reuben Cheruiyot, Odhiambo Romanus Otieno, and George O. Orwa. A boundary corrected non-parametric regression estimator for finite population total. International Journal of Statistics and Probability, 8 (3):83–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/39229.

E:2019:CEG

[435] Sandhya E and Latha C. M. Compound extended geometric distribution and some of its properties. International Journal of Statistics and Probability, 8(3):96–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/39284.

Smith:2019:RAIc

[436] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 8, no. 3. International Journal of Statistics and Probability, 8(3):114–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/39322.

Chukwudum:2019:RPL

[437] Queensley C. Chukwudum. Reinsurance pricing of large motor insurance claims in Nigeria: an extreme value analysis. International Journal of Statistics and Probability, 8(4):1–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/39750.

Megerdichian:2019:EAO

[438] Aren Megerdichian. Economic analysis of optimal sample size in legal proceedings. International Journal of Statistics and Probability, 8 (4):13–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/39758.

Hamedani:2019:CFR

[439] G. G. Hamedani. On characterizations of four recently introduced distributions: Two continuous and two discrete. International Journal of Statistics and Probability, 8(4):25–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/39778.

Sifriyani:2019:SHT

[440] Sifriyani Sifriyani. Simultaneous hypothesis testing of multivariable nonparametric spline regression in the GWR model. International Journal of Statistics and Probability, 8(4):32–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/39865.

Furterer:2019:IHQ

[441] Sandra L. Furterer and Ethling Hernandez. Improving the healthcare quality measurement system using attribute agreement analysis assessing the presence and stage of pressure ulcers. International Journal of Statistics and Probability, 8(4):47–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/39866.

Merie:2019:MID

[442] Ahmed Merie and Myron Hlynka. Medical intervention for disease stages using game theory, Markov chains, and Bayesian inference. International Journal of Statistics and Probability, 8(4):60–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/39880.

Pacifico:2019:ICM

[443] Antonio Pacifico. International co-movements and business cycles synchronization across advanced economies: a SPBVAR evidence. International Journal of Statistics and Probability, 8(4):68–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/39986.

Onwuegbuche:2019:AEV

[444] Faithful C. Onwuegbuche, Alpha B. Kenyatta, Steeven B. Affognon, Exavery P. Enock, and Mary O. Akinade. Application of extreme value theory in predicting climate change induced extreme rainfall in Kenya. International Journal of Statistics and Probability, 8(4):85–??, ???? 2019. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/40020.

Smith:2019:RAId

[445] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 8, no. 4. International Journal of Statistics and Probability, 8(4):95–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/40058.

Schweizer:2019:CCD

[446] Karl Schweizer and Siegbert Reiß. On the contextual conditions driving a difficulty factor. International Journal of Statistics and Probability, 8(5):1–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/40272.

Jones:2019:AGM

[447] Riley Jones and Adriana Ocejo. Assessing guaranteed minimum income benefits and rationality of exercising reset options in variable annuities. International Journal of Statistics and Probability, 8(5): 13–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/40382.

Laverty:2019:VSI

[448] William. H. Laverty and Ivan. W. Kelly. Visualizing state identification in Auto-Regressive Hidden Markov (ARHMM) models with the forward and backward algorithms using Excel. International Journal of Statistics and Probability, 8(5):25–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/40383.

KA:2019:ENH

[449] Diouma Sira KA, George Otieno Orwa, and Oscar Ngesa. Exponentiated Nadarajah–Haghighi–Poisson distribution. International Journal of Statistics and Probability, 8(5):34–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/40406.

Omair:2019:LFT

[450] Maha A. Omair, Abdullah A. Al-Shiha, and Ruba A. Alyafi. Lack-of-fit testing for polynomial regression models without replications. International Journal of Statistics and Probability, 8(5):49–??, ???? 2019. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/40416.

Ando:2019:BIV

[451] Shuji Ando. A bivariate index for visually measuring marginal inhomogeneity in square tables. International Journal of Statistics and Probability, 8(5):58–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/40417.

Nandram:2019:BAS

[452] Balgobin Nandram and Yuan Yu. Bayesian analysis of sparse counts obtained from the unrelated question design. International Journal of Statistics and Probability, 8(5):66–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/40449.

J:2019:GGD

[453] Kiche J, Oscar Ngesa, and George Orwa. On generalized gamma distribution and its application to survival data. International Journal of Statistics and Probability, 8(5):85–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/40527.

Smith:2019:RAIe

[454] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 8, no. 5. International Journal of Statistics and Probability, 8(5):103–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/40530.

Chipepa:2019:NGF

[455] Fastel Chipepa, Broderick O. Oluyede, and Boikanyo Makubate. A new generalized family of odd Lindley-G distributions with application. International Journal of Statistics and Probability, 8(6):1–??, ???? 2019. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/40838.

Osho:2019:GFT

[456] G. Solomon Osho. A general framework for time series forecasting model using autoregressive integrated moving average-ARIMA and transfer functions. International Journal of Statistics and Probability, 8(6): 23–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/40839.

Huillet:2019:PPA

[457] Thierry E. Huillet. Partitioning problems arising from independent shifted-geometric and exponential samples with unequal intensities. International Journal of Statistics and Probability, 8(6):31–??, ???? 2019. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/40996.

Zhang:2019:FPC

[458] Yiping Zhang, Myron Hlynka, and Percy H. Brill. First passage and collective marks. International Journal of Statistics and Probability, 8 (6):47–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/40997.

Alzaghal:2019:NFG

[459] Ahmad Alzaghal and Duha Hamed. New families of generalized Lomax distributions: Properties and applications. International Journal of Statistics and Probability, 8(6):51–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/41029.

Li:2019:LES

[460] Jing Li and Xueyan Li. Liu estimator in semiparametric partially linear varying coefficient models. International Journal of Statistics and Probability, 8(6):69–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/41294.

Park:2019:RMA

[461] Jason Hong Jae Park. Random measure algebras under o-dot product and Morse-Transue integral convolution. International Journal of Statistics and Probability, 8(6):73–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/41303.

Hare:2019:EWP

[462] David R. Hare and Faisal Kaleem. Expected winning probabilities in sequential truels under uniform distributions. International Journal of Statistics and Probability, 8(6):82–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/41315.

Stephen:2019:SPR

[463] Adebowale Ayo Stephen, Asa Soladoye, Abe John Olugbenga, and Oyinlola Funmilola Folasade. Sex preference, religion and ethnicity roles in fertility among women of childbearing age in Nigeria: Examining the links using zero-inflated Poisson regression model. International Journal of Statistics and Probability, 8(6):91–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/41321.

Smith:2019:RAIf

[464] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 8, no. 6. International Journal of Statistics and Probability, 8(6):107–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/41324.

Kawasaki:2020:TTT

[465] Tamae Kawasaki, Toshiki Naito, and Takashi Seo. T^{2} type test statistic and simultaneous confidence intervals for two sub-mean vectors. International Journal of Statistics and Probability, 9(1):1–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/41584.

Pimenta:2020:APC

[466] Cristie Diego Pimenta, Messias Borges Silva, Rose Lima de Morais Campos, and Fabrício Maciel Gomes. Application of principal component analysis and response surface methodology in the process of steel wire tempering. International Journal of Statistics and Probability, 9(1): 9–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/41641.

Bar-Lev:2020:ITI

[467] Shaul K. Bar-Lev. Independent, tough identical results: The class of Tweedie on power variance functions and the class of Bar-Lev and Enis on reproducible natural exponential families. International Journal of Statistics and Probability, 9(1):30–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/41652.

Alkubaisi:2020:REL

[468] Ghaith Abdulsattar A. Jabbar Alkubaisi. The role of ensemble learning in stock market classification model accuracy enhancement based on naive Bayes classifiers. International Journal of Statistics and Probability, 9 (1):36–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/41680.

Haq:2020:MOP

[469] Muhammad Ahsan Ul Haq, G. G. Hamedani, M. Elgarhy, and Pedro Luiz Ramos. Marshall–Olkin power Lomax distribution: Properties and estimation based on complete and censored samples. International Journal of Statistics and Probability, 9(1):48–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/41684.

Patterson:2020:DMM

[470] Joshua Patterson and Andrzej Korzeniowski. Decomposition of M/M/1 with unreliable service and a working vacation. International Journal of Statistics and Probability, 9(1):63–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/41735.

Nanthakumar:2020:CAC

[471] A. Nanthakumar. A comparison of Archimedean copula models for approximating bivariate skew-normal distribution. International Journal of Statistics and Probability, 9(1):70–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/41740.

Smith:2020:RAIa

[472] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 9, no. 1. International Journal of Statistics and Probability, 9(1):87–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/41741.

Wei:2020:EEI

[473] Chuanhua Wei, Nana Zheng, and Ke Tian. Efficient estimation of intervalvalued symbolic data regression model. International Journal of Statistics and Probability, 9(2):1–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/42082.

BODUNWA:2020:ODL

[474] O. K. BODUNWA and O. A. FASORANBAKU. D-optimal design in linear model with different heteroscedasticity structures. International Journal of Statistics and Probability, 9(2):7–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/42132.

Mode:2020:AAM

[475] Charles J. Mode. An algorithmic approach to modelling the co-evolution of parasites and their hosts. International Journal of Statistics and Probability, 9(2):13–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/42133.

Kungu:2020:OSD

[476] Ngigi Peter Kung’u, J. K. Arap Koske, and Josphat K. Kinyanjui. Doptimal slope design for second degree Kronecker model mixture experiment with three ingredients. International Journal of Statistics and Probability, 9(2):30–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/42157.

Kinyanjui:2020:BEP

[477] Josphat. K. Kinyanjui and Betty. C. Korir. Bayesian estimation of parameters of Weibull distribution using LINEX error loss function. International Journal of Statistics and Probability, 9(2):38–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/42166.

Smith:2020:RAIb

[478] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 9, no. 2. International Journal of Statistics and Probability, 9(2):54–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/42167.

Kim:2020:PBA

[479] Steven B. Kim and Jeffrey O. Wand. A paradox in Bland–Altman analysis and a Bernoulli approach. International Journal of Statistics and Probability, 9(3):1–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/42352.

Obasohan:2020:CWM

[480] Phillips Edomwonyi Obasohan. Comparing weighted Markov chain and auto-regressive integrated moving average in the prediction of under-5 mortality annual closing rates in Nigeria. International Journal of Statistics and Probability, 9(3):13–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/42354.

Khorsheed:2020:IWP

[481] Eman Khorsheed. The impact of women parliamentarians on economic growth: Modelling & statistical analysis of empirical global data. International Journal of Statistics and Probability, 9(3):23–??, ???? 2020. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/42498.

Hamedani:2020:CEG

[482] Gholamhossein G. Hamedani and Shirin Nezampour. Characterizations of exponentiated generalized power Lindley and geometric-zero truncated Poisson distributions. International Journal of Statistics and Probability, 9(3):36–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/42512.

Bacilio:2020:GMP

[483] Guerrero-Escamilla Juan Bacilio, Lopez-Perez Socrates, and Jiménez-Proa Erik Gerardo. Gaussian model to predict the risk of developing type 2 diabetes mellitus in Mexican population taking as a reference risk factors. International Journal of Statistics and Probability, 9 (3):42–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/42529.

Marzjarani:2020:CGL

[484] Morteza Marzjarani. A comparison of a general linear model and the ratio estimator. International Journal of Statistics and Probability, 9 (3):54–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/42530.

Almutairi:2020:CNG

[485] Abdulwahab Almutairi. Combined nearest greedy algorithm with randomized iterated greedy algorithm to solve waste collection problem. International Journal of Statistics and Probability, 9(3):66–??, ???? 2020. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/42581.

Smith:2020:RAIc

[486] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 9, no. 3. International Journal of Statistics and Probability, 9(3):72–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/42618.

Andrei:2020:BAA

[487] Mihnea S. Andrei and John S. J. Hsu. A Bayesian approach for asset allocation. International Journal of Statistics and Probability, 9 (4):1–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/42740.

Chipepa:2020:TLM

[488] Fastel Chipepa, Broderick Oluyede, and Boikanyo Makubate. The Topp– Leone Marshall–Olkin-G family of distributions with applications. International Journal of Statistics and Probability, 9(4):15–??, ???? 2020. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/42741.

Tengaa:2020:SMH

[489] Peter Emeda Tengaa, Samuel Mwalili, and George Otieno Orwa. Stochastic modeling for HIV and AIDS epidemics with viral load detectability. International Journal of Statistics and Probability, 9(4): 33–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/42999.

He:2020:SIA

[490] Hua He, Xuenan Mi, Jerry Cornell, Wan Tang, Tanika Kelly, Hui Shen, Hongwen Deng, and Yan Du. Statistical issues on analysis of censored data due to detection limit. International Journal of Statistics and Probability, 9(4):49–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/43016.

Beauducel:2020:OLI

[491] André Beauducel and Norbert Hilger. Overcoming limitations of the independent clusters model for CFA by means of Bayes-estimation and buffered simple structure. International Journal of Statistics and Probability, 9(4):62–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/43046.

Sharma:2020:DHP

[492] Richa Sharma and Ajay Pandey. Does higher percentages of women with higher education within district impacts individuals use of contraception in Uttar Pradesh? International Journal of Statistics and Probability, 9(4):78–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/43155.

Smith:2020:RAId

[493] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 9, no. 4. International Journal of Statistics and Probability, 9(4):85–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/43162.

Almazah:2020:EPE

[494] Mohammed Mohammed Ahmed Almazah. Estimation of the parameters for the exponentiated Weibull distribution based on progressive Type-II censoring scheme. International Journal of Statistics and Probability, 9(5):1–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/43344.

Hemeda:2020:SIE

[495] Saeed E. Hemeda and Ali M. Abdallah. Sinh inverted exponential distribution: Simulation & application to neck cancer disease. International Journal of Statistics and Probability, 9(5):11–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/43345.

Liu:2020:RTI

[496] Dan Liu, Yuan Du, Yasin Khadem Charvadeh, Jingyu Cui, Li-Pang Chen, Gansen Deng, Qihuang Zhang, Kaida Cai, Joy He, Wenqing He, and Grace Y. Yi. A real time and interactive Web-based platform for visualizing and analyzing COVID-19 in Canada. International Journal of Statistics and Probability, 9(5):23–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/43346.

Islam:2020:NTT

[497] Khairul Islam and Tanweer J. Shapla. A new transformed t-test for skewed data: a goodness-of-fit approach. International Journal of Statistics and Probability, 9(5):30–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/43347.

Lim:2020:ESC

[498] Eunji Lim and Kihwan Kim. Estimating smooth and convex functions. International Journal of Statistics and Probability, 9(5):40– ??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/43351.

Escamilla:2020:LMP

[499] Juan Bacilio Guerrero Escamilla, Sócrates López Pérez, and Sonia Bass Zavala. Logistic model to predict the contagion of COVID-19 in Mexico. International Journal of Statistics and Probability, 9(5):49– ??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/43440.

Azumah:2020:FTS

[500] Karim Azumah, Ananda Omutokoh Kube, and Bashiru Imoro Ibn Saeed. Functional time series analysis of land surface temperature. International Journal of Statistics and Probability, 9(5):61–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/43556.

Obuya:2020:RWD

[501] Emilly A. Obuya, Prakash C. Joshi, Thomas A. Gray, Thomas C. Keane, and Wayne E. Jones Jr. Revisit the Wishart distributionm. International Journal of Statistics and Probability, 9(5):79–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/43557.

Kalampakas:2020:SAB

[502] Argyrios Kalampakas and Georgios C. Makris. Statistical analysis to bitcoin transactions network. International Journal of Statistics and Probability, 9(5):85–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/43581.

Smith:2020:RAIe

[503] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 9, no. 5. International Journal of Statistics and Probability, 9(5):97–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/43624.

Bidounga:2020:NBC

[504] Rufin Bidounga, Evrand Giles Brunel Mandangui Maloumbi, Réolie Foxie Mizélé Kitoti, and Dominique Mizère. The new bivariate Conway– Maxwell–Poisson distribution obtained by the crossing method. International Journal of Statistics and Probability, 9(6):1–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/43714.

Kim:2020:CIC

[505] Dong yun Kim and Yanhong Wu. Confidence interval for change point in hazard rate with staggered entry. International Journal of Statistics and Probability, 9(6):9–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/43715.

Mallipaddi:2020:EST

[506] Venkata Mallipaddi and Michael Anderson. Evaluating safety of Type-A weaving sections using geometric and traffic operational factors. International Journal of Statistics and Probability, 9(6):21–??, ???? 2020. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/43798.

Ogura:2020:ESP

[507] Toru Ogura, Takatoshi Sugiyama, and Nariaki Sugiura. Estimation of the shape parameter of a wear-out failure period for a three-parameter Weibull distribution in a small sample. International Journal of Statistics and Probability, 9(6):39–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/43799.

Agrawal:2020:ISP

[508] Yogesh Agrawal, Faisal Kaleem, and David Hare. An introduction to sequential pentuals. International Journal of Statistics and Probability, 9(6):49–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/43853.

Onozawa:2020:NTS

[509] Mizuki Onozawa, Ayaka Yagi, and Takashi Seo. New test statistics for one and two mean vectors with two-step monotone missing data. International Journal of Statistics and Probability, 9(6):56–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/43943.

Konane:2020:SMF

[510] Victorien F. Konane and Ali Traore. Statistical modeling and forecast of the corona-virus disease (COVID-19) in Burkina Faso. International Journal of Statistics and Probability, 9(6):76–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/43967.

Ogunde:2020:PIA

[511] A. A. Ogunde, S. T. Fayose, B. Ajayi, and D. O. Omosigho. Properties, inference and applications of alpha power extended inverted Weibull distribution. International Journal of Statistics and Probability, 9(6): 90–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/43989.

Young:2020:TEC

[512] Phil D. Young, Joshua D. Patrick, and Dean M. Young. Two explicit characterizations of the general nonnegative-definite covariance matrix structure for equality of BLUEs, WLSEs, and LSEs. International Journal of Statistics and Probability, 9(6):108–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/43990.

Hapsari:2020:EVP

[513] Meilina Retno Hapsari, Suci Astutik, and Loekito Adi Soehono. Estimation of VECM parameter using Bayesian approach: an application to analysis of macroeconomic variables. International Journal of Statistics and Probability, 9(6):113–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/44018.

Smith:2020:RAIf

[514] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 9, no. 6. International Journal of Statistics and Probability, 9(6):124–??, ???? 2020. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/44048.

Sawadogo:2021:CCT

[515] Amadou Sawadogo, Dominique Lafon, and Simplice Dossou-Gbété. On the classification of colored textures from a texture-ranking experiment: Observers ability of discrimination quantification. International Journal of Statistics and Probability, 10(1):1–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/44327.

Tenenholtz:2021:SMR

[516] John Tenenholtz, Florence George, and Sneh Gulati. Some multiple regression models for the number of COVID-19 cases and deaths in the United States. International Journal of Statistics and Probability, 10 (1):28–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44328.

Nakamura:2021:CST

[517] Tomoshige Nakamura and Mihoko Minami. Causal subclassification tree algorithm and robust causal effect estimation via subclassification. International Journal of Statistics and Probability, 10(1):40–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/44329.

Andrei:2021:BAL

[518] Mihnea S. Andrei and John S. J. Hsu. A Bayesian approach for large asset allocation. International Journal of Statistics and Probability, 10 (1):58–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44345.

Lim:2021:CPC

[519] Eunji Lim. Consistency of penalized convex regression. International Journal of Statistics and Probability, 10(1):69–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/44423.

Kellner:2021:DCD

[520] Caroline Kellner and Lieven Nils Kennes. Decision criteria during joint modelling of efficacy and safety with MCP-Mod. International Journal of Statistics and Probability, 10(1):79–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/44424.

Gaafar:2021:EET

[521] Mohammad Ibrahim Ahmmad Soliman Gaafar. An empirical evaluation of a test procedure for the median of symmetrical and asymmetrical populations using an interpolated nonparametric confidence interval. International Journal of Statistics and Probability, 10(1):85–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/44479.

Smith:2021:RAIa

[522] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 10, no. 1. International Journal of Statistics and Probability, 10(1):106–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/44494.

Andrei:2021:DCM

[523] Mihnea S. Andrei, Sujit K. Ghosh, and Jian Zou. Dynamic correlation multivariate stochastic volatility black-litterman with latent factors. International Journal of Statistics and Probability, 10(2):1–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/44571.

Koech:2021:ERO

[524] Ben Kiprono Koech. Estimation of receiver operating characteristic surface using mixtures of finite Polya trees (MFPT). International Journal of Statistics and Probability, 10(2):18–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/44641.

Song:2021:AGS

[525] Ge Song, Jiahui Yuan, and Charlie Cheng-Jie Ji. Application of Gibbs sampling in modelling the utilization rate of raw materials for drug coating. International Journal of Statistics and Probability, 10 (2):28–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44642.

Kungu:2021:OSD

[526] Ngigi Peter Kung’u, J. K. Arap Koske, and Josphat K. Kinyanjui. A-optimal slope design for second degree Kronecker model mixture experiment with four ingredients with application in selected fruits blending. International Journal of Statistics and Probability, 10(2): 36–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44694.

Magrini:2021:EDB

[527] Alessandro Magrini. Efficient decomposition of Bayesian networks with non-graded variables. International Journal of Statistics and Probability, 10(2):52–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44695.

Escamilla:2021:CST

[528] Juan Bacilio Guerrero Escamilla and Arquímedes Avilés Vargas. Crosssectional and time series data as the basis for panel modelling: The case of kidnappings in México from 2010 to 2019. International Journal of Statistics and Probability, 10(2):68–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/44783.

Choi:2021:LSP

[529] Jai Won Choi and Balgobin Nandram. Large sample problems. International Journal of Statistics and Probability, 10(2):81–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/44784.

Mahmoud:2021:PVS

[530] Hamdy F. F. Mahmoud. Parametric versus semi and nonparametric regression models. International Journal of Statistics and Probability, 10 (2):90–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44804.

Lopez-Rios:2021:LAC

[531] Ruy Lopez-Rios. A Q-learning approach to a consumption–investment problem. International Journal of Statistics and Probability, 10(2): 110–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44805.

Makubate:2021:MOH

[532] Boikanyo Makubate, Fastel Chipepa, Broderick Oluyede, and Peter O. Peter. The Marshall–Olkin half logistic-g family of distributions with applications. International Journal of Statistics and Probability, 10(2): 120–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44808.

Abdul-Aziz:2021:CPE

[533] A. R. Abdul-Aziz, Albert Luguterah, and Bashiru I. I. Saeed. Comparative performance of estimation maximization among residual estimators: a structural equation modelling perspective. International Journal of Statistics and Probability, 10(2):138–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/44815.

Smith:2021:RAIb

[534] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 10, no. 2. International Journal of Statistics and Probability, 10(2):146–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/44819.

Wei:2021:ELR

[535] Chuanhua Wei and Xiaoxiao Ma. Empirical likelihood ratio test for seemingly unrelated regression models. International Journal of Statistics and Probability, 10(3):1–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/44862.

Ogunde:2021:LTI

[536] Adebisi Ade Ogunde, Gbenga Adelekan Olalude, Oyebimpe Emmanuel Adeniji, and Kayode Balogun. Lehmann Type II Frechet Poisson distribution: Properties, inference and applications as a life time distribution. International Journal of Statistics and Probability, 10(3): 8–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44974.

Merino:2021:ACP

[537] Lucila Muniz Merino, Bulmaro Juarez Hernandez, and Hugo A. Cruz Suarez. Analysis of change points with Bayes factor, thresholds, and CUSUM. International Journal of Statistics and Probability, 10(3): 31–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/44975.

Arshad:2021:SPN

[538] Muhammad Z. Arshad, Muhammad Z. Iqbal, and Alya Al Mutairi. Statistical properties of a new bathtub shaped failure rate model with applications in survival and failure rate data. International Journal of Statistics and Probability, 10(3):49–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/44976.

Qin:2021:SLI

[539] Lu Qin, Jihong Zhang, Xinya Liang, and Qianqian Pan. R-squared of a latent interaction in structural equation model: a tutorial of using R. International Journal of Statistics and Probability, 10(3):69–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http:/

/www.ccsenet.org/journal/index.php/ijsp/article/view/0/45113.

Shi:2021:GMF

[540] Qun Shi. Generalized mean-field fractional BSDEs with non-Lipschitz coefficients. International Journal of Statistics and Probability, 10(3):

77–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-

7040 (electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/45114.

Muse:2021:LLD

[541] Abdisalam Hassan Muse, Samuel M. Mwalili, and Oscar Ngesa. On the log-logistic distribution and its generalizations: a survey. International Journal of Statistics and Probability, 10(3):93–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.

ccsenet.org/journal/index.php/ijsp/article/view/0/45138.

Kelner:2021:CAC

[542] Moshe Kelner, Zinoviy Landsman, and Udi E. Makov. Compound Archimedean copulas. International Journal of Statistics and Probability, 10(3):126–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/45156.

Lin:2021:MPMa

[543] Hejie Lin and Tsung-Wu Lin. Mechanical proof of the Maxwell–Boltzmann speed distribution with analytical integration. International Journal of Statistics and Probability, 10(3):135–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/45187.

Eniade:2021:EEA

[544] Olanrewaju Davies Eniade, Joshua Odunayo Akinyemi, Oyindamola Bidemi Yusuf, Rotimi Felix. Afolabi, and Olufunmilayo I. Fawole. Effect of education on attitude towards domestic violence in Nigeria: an exploration using propensity score methodology. International Journal of Statistics and Probability, 10(3):154–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.org/

journal/index.php/ijsp/article/view/0/45205.

Smith:2021:RAIc

[545] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 10, no. 3. International Journal of Statistics and Probability, 10(3):164–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/0/45206.

Nanthakumar:2021:DTBa

[546] A. Nanthakumar. A diagnostic test based on a 9-component mixture Gaussian copula model. International Journal of Statistics and Probability, 10(4):1–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL http://www.ccsenet.org/journal/index.
php/ijsp/article/view/0/45299.

Nanthakumar:2021:DTBb

[547] A. Nanthakumar. A diagnostic test based on a 9-component mixture Gaussian copula model. International Journal of Statistics and Probability, 10(4):1–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/45299.

Nzokem:2021:SEM

[548] A. H. Nzokem. SIS epidemic model birth-and-death Markov chain approach. International Journal of Statistics and Probability, 10(4):10–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/45361.

Lin:2021:MPMb

[549] Hejie Lin and Tsung-Wu Lin. Mechanical proof of the Maxwell–Boltzmann speed distribution with numerical iterations. International Journal of Statistics and Probability, 10(4):21–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/45427.

Makubate:2021:NLB

[550] Boikanyo Makubate, Broderick Oluyede, and Morongwa Gabanakgosi. A new Lindley-Burr XII distribution: Model, properties and applications. International Journal of Statistics and Probability, 10(4):33–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/45408.

Tasnim:2021:USR

[551] Sanjida Tasnim. Use of shape restricted regression methods for fitting model of per capita GDP: a global economic scenario of 2018. International Journal of Statistics and Probability, 10(4):52–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://

ccsenet.org/journal/index.php/ijsp/article/view/0/45409.

Ouedraogo:2021:MML

[552] Ouindllassida Jean-Etienne Ou’edraogo, Edoh Katchekpele, and Simplice Dossou-Gb’et’e. Marginalized maximum likelihood for parameters estimation of the three parameter Weibull distribution. International Journal of Statistics and Probability, 10(4):62–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/45441.

Ogunde:2021:HEP

[553] Adebisi Ade Ogunde, Victoria Eshomomoh Laoye, Ogbonnaya Nzie Ezichi, and Kayode Oguntuase Balogun. Harris extended power Lomax distribution: Properties, inference and applications. International Journal of Statistics and Probability, 10(4):77–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.

org/journal/index.php/ijsp/article/view/0/45482.

Pimenta:2021:ASM

[554] Cristie Diego Pimenta, Messias Borges Silva, Fernando Augusto Silva Marins, and Aneirson Francisco da Silva. Application of statistical monitoring using autocorrelated data and with the influence of multicollinearity in a steel process. International Journal of Statistics and Probability, 10(4):96–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/45489.

Aga:2021:EEW

[555] Mosisa Aga. Edgeworth expansion for the Whittle maximum likelihood estimator of linear regression processes with long memory residuals. International Journal of Statistics and Probability, 10(4):119–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/45494.

Islam:2021:ADR

[556] Khairul Islam and Tanweer J. Shapla. Application of discrete regression models for analyzing K–8 students nonchronic absenteeism in the United States. International Journal of Statistics and Probability, 10(4):127–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/45495.

Magrini:2021:SFM

[557] Alessandro Magrini. A stochastic frontier model to assess agricultural ecoefficiency of European countries in 1990–2019. International Journal of Statistics and Probability, 10(4):138–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/45496.

Louzayadio:2021:WPD

[558] Chedly Gelin Louzayadio, Rodnellin Onesime Malouata, and Michel Diafouka Koukouatikissa. A weighted Poisson distribution for underdispersed count data. International Journal of Statistics and Probability, 10(4):157–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/45530.

Hertz:2021:UAP

[559] David Hertz. Unified approach to probability problems and estimation algorithms associated with symmetric functions. International Journal of Statistics and Probability, 10(4):166–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/45534.

Smith:2021:RAId

[560] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 10, no. 4. International Journal of Statistics and Probability, 10(4):183–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/45535.

Muiftah:2021:EPT

[561] Mohamed S. A. Muiftah and Samir K. Ashour. Estimation of the parameters of type-II discrete Weibull distribution under type-I censoring. International Journal of Statistics and Probability, 10(5):1–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/45791.

Kelner:2021:FCA

[562] Moshe Kelner, Zinoviy Landsman, and Udi E. Makov. Fitting compound Archimedean copulas to data for modeling electricity demand. International Journal of Statistics and Probability, 10(5):20–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/45792.

Hertz:2021:IHL

[563] David Hertz. Improved Hoeffding’s lemma and Hoeffding’s tail bounds. International Journal of Statistics and Probability, 10(5):27–??, ????

2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/45793.

Lim:2021:SBO

[564] Eunji Lim. Simulation-based optimization for convex functions over discrete sets. International Journal of Statistics and Probability, 10(5):31–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/45794.

Chen:2021:BBC

[565] Wei Chen and Fengling Ren. Bootstrap based confidence interval estimation of quantiles for current status data. International Journal of Statistics and Probability, 10(5):38–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/45816.

Davydenko:2021:APF

[566] Andrey Davydenko and Paul Goodwin. Assessing point forecast bias across multiple time series: Measures and visual tools. International Journal of Statistics and Probability, 10(5):46–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.

org/journal/index.php/ijsp/article/view/0/45817.

Mbunzi:2021:SIL

[567] Stephen M. Mbunzi, Joseph K. Mungatu, Anthony G. Waititu, Samuel M. Mwalili, Kenneth O. Ogila, Thomas N. O. Achia, and Daniel Nthiwa. Severity of infestation levels of Tunga Penetrans in Central Kenya: a Bayesian cumulative logit model. International Journal of Statistics and Probability, 10(5):70–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/45819.

Wang:2021:CWT

[568] Xiaoming Wang, Sukun Wang, and Warren Kindzierski. Control of weekly time trend in time-stratified case-crossover design. International Journal of Statistics and Probability, 10(5):85–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/45843.

Smith:2021:RAIe

[569] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 10, no. 5. International Journal of Statistics and Probability, 10(5):98–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/45866.

Mason:2021:CEM

[570] Zadan P. H. Mason and Mahrokh Najaf. Characterizations of the exponentiated Marshall–Olkin discrete uniform distribution. International Journal of Statistics and Probability, 10(6):1–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.

org/journal/index.php/ijsp/article/view/0/46011.

Nandram:2021:INP

[571] Balgobin Nandram, Jai Won Choi, and Yang Liu. Integration of nonprobability and probability samples via survey weights. International Journal of Statistics and Probability, 10(6):5–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/46130.

Aldossary:2021:UGC

[572] Fahad Aldossary, Tamer Abouelmagd, and A. A. E. Ahmed. A unified general class of interdistributional income inequality measures based on weighted incomplete moments. International Journal of Statistics and Probability, 10(6):22–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/46132.

Al-Aqtash:2021:GBX

[573] Raid Al-Aqtash, Avishek Mallick, G. G. Hamedani, and Mahmoud Aldeni. On the Gumbel–Burr XII distribution: Regression and application. International Journal of Statistics and Probability, 10(6):31–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/46173.

Mafale:2021:MPM

[574] Ndubano Mafale, Dismas Ntirampeba, and Jacob Ong’ala. Multidimensional poverty modeling for Namibia using the beta distribution. International Journal of Statistics and Probability, 10(6):47–??, ???? 2021. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/46201.

Badyalina:2021:MGM

[575] Basri Badyalina, Ani Shabri, Nurkhairany Amyra Mokhtar, Mohamad Faizal Ramli, Muhammad Majid, and Muhammad Yassar Yusri. Modified group method of data handling for flood quantile prediction at ungauged site. International Journal of Statistics and Probability, 10(6): 57–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040

(electronic). URL https://ccsenet.org/journal/index.php/ijsp/
article/view/0/46202.

Smith:2021:RAIf

[576] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 10, no. 6. International Journal of Statistics and Probability, 10(6):67–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/46203.

Nuga:2022:OSP

[577] Oluwole A. Nuga, Abba Zakirai Abdulhamid, and Shobanke Emmanuel Omobola Kayode. D-optimal split-plot designs with random whole plot factor and fixed sub-plot factor. International Journal of Statistics and Probability, 11(1):1–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/46493.

Gbenga:2022:APE

[578] Jemilohun Vincent Gbenga and Ipinyomi Reuben Adeyemi. Alpha power extended inverse Weibull Poisson distribution: Properties, inference, and applications to lifetime data. International Journal of Statistics and Probability, 11(1):10–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/46494.

Jiang:2022:MCC

[579] Yu Hang Jiang, Tong Liu, Zhiya Lou, Jeffrey S. Rosenthal, Shanshan Shangguan, Fei Wang, and Zixuan Wu. Markov chain confidence intervals and biases. International Journal of Statistics and Probability, 11(1):29–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/46495.

Sugiyama:2022:PEW

[580] Takatoshi Sugiyama and Toru Ogura. Parameters estimation for wearout failure period of three-parameter Weibull distribution. International Journal of Statistics and Probability, 11(1):40–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://

ccsenet.org/journal/index.php/ijsp/article/view/0/46517.

Pandey:2022:AFB

[581] Ajay Pandey and Richa Sharma. Age at first birth in uttar pradesh: How much has it actually changed? Over inter-NFHS period? International Journal of Statistics and Probability, 11(1):54–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://

ccsenet.org/journal/index.php/ijsp/article/view/0/46534.

Smith:2022:RAIa

[582] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 11, no. 1. International Journal of Statistics and Probability, 11(1):67–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/46535.

Harripersaud:2022:DCP

[583] Ashwannie Harripersaud. Demystifying the concept of probability. International Journal of Statistics and Probability, 11(2):1–??, ???? 2022. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/46744.

Gbenga:2022:APP

[584] Jemilohun Vincent Gbenga and Ipinyomi Reuben Adeyemi. Alpha power Poisson-G distribution with an application to Burr II distribution lifetime data. International Journal of Statistics and Probability, 11(2):8–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/46745.

AlMutairi:2022:NOF

[585] Alya Al Mutairi and Muhammad Z. Arshad. A new odd Fréchet Lehmann type II-G family of distributions: a power function distribution with theory and applications. International Journal of Statistics and Probability, 11(2):29–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/46746.

Wilcox:2022:IAQ

[586] Rand R. Wilcox. Inferences about a quantile shift measure of effect size when there is a covariate. International Journal of Statistics and Probability, 11(2):52–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/46782.

Young:2022:RMA

[587] S. Stanley Young, Kai-Chieh Cheng, Jin Hua Chen, Shu-Chuan Chen, and Warren B. Kindzierski. Reliability of a meta-analysis of air qualityasthma cohort studies. International Journal of Statistics and Probability, 11(2):61–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/46815.

Smith:2022:RAIb

[588] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 11, no. 2. International Journal of Statistics and Probability, 11(2):77–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/46832.

Ogura:2022:CVW

[589] Toru Ogura and Chihiro Shiraishi. Cutoff value for Wilcoxon–Mann– Whitney test by minimum P -value: Application to COVID-19 data. International Journal of Statistics and Probability, 11(3):1–??, ???? 2022. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/46876.

Huang:2022:BBC

[590] Jie Huang, Haiming Zhou, and Nader Ebrahimi. Bayesian bivariate cure rate models using copula functions. International Journal of Statistics and Probability, 11(3):9–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/46877.

Sarpong:2022:IBW

[591] Smart Asomaning Sarpong and Akwasi Agyei. Interactions between water level, crude oil, and hydroelectric power generation in Ghana; a structured vector auto regression approach. International Journal of Statistics and Probability, 11(3):22–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/47099.

Al-Saleh:2022:ABP

[592] Mohammad Fraiwan Al-Saleh and Doaa Suhail Obeidat. Approximation of the binomial probability function using the discrete normal distribution. International Journal of Statistics and Probability, 11(3):32–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/47100.

Young:2022:SRD

[593] S. Stanley Young and Warren B. Kindzierski. Statistical reliability of a diet-disease association meta-analysis. International Journal of Statistics and Probability, 11(3):40–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/47119.

Smith:2022:RAIc

[594] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 11, no. 3. International Journal of Statistics and Probability, 11(3):51–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/47126.

Angko:2022:DEF

[595] William Angko and Joseph Kwame Wulifan. Demographic effects on fertility among reproductive aged women in Ghana. a demographic and health survey analysis. International Journal of Statistics and Probability, 11(4):1–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/47398.

Nanthakumar:2022:CVH

[596] A. Nanthakumar. A comparison of vine and hierarchical copulas as discriminants. International Journal of Statistics and Probability, 11(4): 13–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040

(electronic). URL https://ccsenet.org/journal/index.php/ijsp/
article/view/0/47407.

Nystrom:2022:URP

[597] Emily Nystrom, Julia L. Sharp, and William C. Bridges, Jr. Using residual plots to distinguish cases of predictor omission in linear models. International Journal of Statistics and Probability, 11(4):25–??, ???? 2022. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/47417.

Nwanyibuife:2022:ZTP

[598] Okechukwu Bridget Nwanyibuife, Okechukwu Angela Unna, and Amahia Godwin Nwanzu. Zero truncated Poisson Harris Weibull distribution: Properties and applications to lifetime datas. International Journal of Statistics and Probability, 11(4):37–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://

ccsenet.org/journal/index.php/ijsp/article/view/0/47418.

Ikpang:2022:MAR

[599] Ikpang Nkereuwem Ikpang and Ekom obong Jackson Okon. Modeling average rainfall in Nigeria with artificial neural network (ANN) models and seasonal autoregressive integrated moving average (SARIMA) models. International Journal of Statistics and Probability, 11(4):53–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/47422.

Smith:2022:RAId

[600] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 11, no. 4. International Journal of Statistics and Probability, 11(4):63–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/47423.

George:2022:ALR

[601] Florence George, Sultana Mubarika Rahman Chowdhury, and Sneh Gulati. Application of logistic regression on heart disease data and a review of some standardization methods. International Journal of Statistics and Probability, 11(5):1–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/47686.

Hassim:2022:SFF

[602] Nor Hidayah Hassim, Basri Badyalina, Nurkhairany Amyra Mokhtar, Muhammad Zulqarnain Hakim Abd Jalal, Nur Diana Binti Zamani, Lee Chang Kerk, Amir Imran Zainoddin, Ahmad Syahmi Ahmad Fadzil, and Nur Fatihah Shaari. In situ flood frequency analysis used for water resource management in Kelantan River Basin. International Journal of Statistics and Probability, 11(5):8–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/47687.

Janeiro:2022:NMM

[603] Vanderly Janeiro, Robson Marcelo Rossi, Terezinha Aparecida Guedes, Ana Beatriz Tozzo Martins, and Lucimary Afonso dos Santos. Nonlinear mixed models applied to ruminal degradability studies. International Journal of Statistics and Probability, 11(5):18–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://

ccsenet.org/journal/index.php/ijsp/article/view/0/47740.

Sarpong:2022:FHG

[604] Smart Asomaning Sarpong and Akwasi Agyei. Forecasting hydropower generation in Ghana using ARIMA models. International Journal of Statistics and Probability, 11(5):30–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/47741.

Iki:2022:PBD

[605] Kiyotaka Iki and Sadao Tomizawa. Parsimonious bivariate T -distribution type symmetry models for square contingency tables. International Journal of Statistics and Probability, 11(5):44–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.

org/journal/index.php/ijsp/article/view/0/47754.

Smith:2022:RAIe

[606] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 11, no. 5. International Journal of Statistics and Probability, 11(5):53–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/47756.

Beauducel:2022:CPM

[607] André Beauducel and Norbert Hilger. Correlation-preserving mean plausible values as a basis for prediction in the context of Bayesian structural equation modeling. International Journal of Statistics and Probability, 11(6):1–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/47905.

Choi:2022:CCP

[608] Jai Won Choi, Balgobin Nandram, and Boseung Choi. Combining correlated p-values from primary data analyses. International Journal of Statistics and Probability, 11(6):12–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/

journal/index.php/ijsp/article/view/0/47906.

Fernando:2022:RCB

[609] Dimuthu Fernando, Mohammed Alqawba, Manar Samad, and Norou Diawara. Review of copula for bivariate distributions of zero-inflated count time series data. International Journal of Statistics and Probability, 11(6):28–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/47926.

Torres:2022:SSS

[610] Lucia V. P. Torres and Juan B. Guerrero Escamilla. Simple sampling for SARS-CoV-2 infection in Hidalgo. International Journal of Statistics and Probability, 11(6):41–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/47927.

Ziedan:2022:BEI

[611] Dalia Ziedan and Ayman Orabi. E-Bayesian estimation for the inverted Topp–Leone distribution based on type-II censored data. International Journal of Statistics and Probability, 11(6):52–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://

ccsenet.org/journal/index.php/ijsp/article/view/0/47951.

Tashkandy:2022:NBG

[612] Yusra A. Tashkandy. Negative binomial and geometric; bivariate and difference distributions. International Journal of Statistics and Probability, 11(6):60–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print),

1927-7040 (electronic). URL https://ccsenet.org/journal/index.
php/ijsp/article/view/0/47952.

Smith:2022:RAIf

[613] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 11, no. 6. International Journal of Statistics and Probability, 11(6):70–??, ???? 2022. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/47953.

Wilcox:2023:RIG

[614] Rand R. Wilcox and Lai Xu. Regression: Identifying good and bad leverage points. International Journal of Statistics and Probability, 12(1):1–??, ???? 2023. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/48193.

Seham:2023:MTL

[615] M. Seham and I. B. Omnia. On the mixture of Topp–Leone-G class and exponentiated-G class of distribution. International Journal of Statistics and Probability, 12(1):9–??, ???? 2023. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/48194.

Li:2023:VSN

[616] Kexuan Li. Variable selection for nonlinear Cox regression model via deep learning. International Journal of Statistics and Probability, 12(1):21–??, ???? 2023. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/48206.

Yang:2023:BPI

[617] Lingli Yang, Balgobin Nandram, and Jai Won Choi. Bayesian predictive inference under nine methods for incorporating survey weights. International Journal of Statistics and Probability, 12(1):33–??, ???? 2023. CO-DEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https:

//ccsenet.org/journal/index.php/ijsp/article/view/0/48223.

Young:2023:RHC

[618] S. Stanley Young and Warren B. Kindzierski. Reproducibility of health claims in meta-analysis studies of COVID quarantine (stay-at-home) orders. International Journal of Statistics and Probability, 12(1):54–??, ???? 2023. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic).

URL https://ccsenet.org/journal/index.php/ijsp/article/view/
0/48224.

Smith:2023:RAI

[619] Wendy Smith. Reviewer acknowledgements for International Journal of Statistics and Probability, vol. 12, no. 1. International Journal of Statistics and Probability, 12(1):66–??, ???? 2023. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL https://ccsenet.org/journal/

index.php/ijsp/article/view/0/48225.